Repair Briefs - Video recorders, TVs, video projectors, DVD players and computer monitors
The following are repair briefs for various
equipment.The infomation is directed
to technically competant repair engineers.Generic terms have
been used to make this info less model specific,eg terms like
replace transistor Q123 have not been used.
The equipment is video recorders, TVs and computer monitors.
I would be interested in finding
any other repair listings on the internet structured as i have
done ie intended to be less model specific. For convenience using search-engines,
use keyword divdevrep to target these files.
There is no point in contacting me about any of the following, the
repair job may have been done 15 years ago .
I cannot clarify or enlarge on any of the following. Mains is 240V, UK
Modern Chinese saying "If it can be mended,It must be mended"
- Chairman Mao Zedong
Should the location of this file change please use the keyword "divdev7" in
a search engine to find it again
With cassette on deck the take up side was held against
the deck but the supply side was not firm so allowing
to rise.This then allowed the bottom flange of the
supply side spool to rub the bottom of the cassette.
So failing to retract tape into cassette going
to STOP mode.Probably due to permanent torsion in the cross
coupling rod that couples cassette motion from one side to the
other. On supply side external to cassette housing removed
large cog and repositioned 1 tooth advanced from previous position.
Half a tooth repositioning would have been better but of course
not possible.It looked increased possibility of cassette jamming
on the housing when in eject position because of skewed action
but it didn't jam once in the access slot
Aiwa FX2500 vcr
Tape jam preceeded by poor rewind
The slip clutch in the swing arm mechanism of 2 cogs
that engages in REW with the main slip clutch was not
slipping enough so changed the internal spring for a weaker one .
Akai VS F30, 1990 vcr
Despite pressing Clock, Transmit on R/C - necessary
after more than 1 minute off mains.
Checked fuseable Rs in ps , ok
and output voltages (later when working in play mode)
It is possible without removing the bottom plate to
inch round the mode switch motor (in the easy sense ) with screw driver
through cassette carrier. M/c would now switch from standby to
on and work but the cassette was not being held down in cassette-in
The rack that drives the cassette motion had jumped a
tooth on the pinion so only one tooth engaged
at cassette-in position when it should be 2.
In cassette-out the mode position was in an invalid
spot, outside of hand-pushing the cassette in to activate ,
and disabled the machine until inched back.
Forced back and made mods to constrain this rack
in tighter mesh with cog.
Cut down a PTFE through chassis feed-thru pin complete
with pin and forced into gap between rack and metal chassis.
Fixed in place with plenty of hot melt glue against metal
and plastic chassis sections.
The erase head does not pivot on this m/c and
forms part of rack constraint but only anchored
with one screw. The other head locator just pushes through
a hole in the chassis. Srewed down this bit with washers padding
out so tightly held and no longer lifting.
Checking motor motion, desoldered leads and powered
from 9V, .1 to .2 amp in free movement.
Akai VS8 vcr.
No transport functions
A small spring attached to the solenoid linkage
mechanism had broken.
Akai VS8 vcr.
Dead machine ,no display,no rf output.
Failure of STK5225 hybrid regulator.Originally replaced
with one from an unchecked VS4 but this too had a
duff STK5225 it transpired.Fault due to over-heating TA7245 drum motor driver,
leading to overheating of the STK 14volt regulator section
Akai VS8 vcr
Intermittant failure while playing followed by tape
jam on trying to remove the cassette.
Small solder bridge on the MBL8243 on the ps board
Akai VS-8 vcr
Random unsynced "Venetian blinds" of anti-phase coulour
in the image in PB.Instability / oscillation around pin 2 ,
the VCO section of the AN6371 and 4.4357MHz,taking
a 10K resistor from pin 13 to ground cured the oscillation .
Akai VS-8 vcr
Tape jam in play mode.
The large tyred pulley that engages with the capstan flywheel
had the plastic boss cracked so allowing the steel spindle to rotate
and act as a slip-clutch.Demount whole jockey-assembly ,remove
pulley and strengthen by gluing with epoxy and expanded metal
reinforcement and deform the end of the steel spindle by squeezing
/ indenting with an old pair of end cutters to produce a key when placing
back into pulley.
Akai VS512 vcr
Failure of drum to get up to speed unless "kick-started" with a finger
Failure of one of the poles of the 3 pole motor driver TA7245F on drum motor board.
Akai VS F10
No cassette loading
The half cog that links the cassette operating metal tooth rack
to the rest of the cassette carrier had a broken tooth.The flimsey
clip that locates this half cog on its shaft and rotates
around a curved slot in the metal casing had failed and then misregistration
had allowed the tooth to break.Replaced with a 20 tooth fibre
reinforced nylon complete cog from some other equipment.Opening out
the plastic pivot piece to this cog and using a bolt and 2 lock nuts to
secure the replacement full cog.
Akai VS G245
Mode switch mechanism out of synch.
Small plastic guide peg obscured from view had sheared off.
This peg is under the main mode change rack that goes the width of
the deck. It restrains the reactive force from the small drive
cog on the main locus cam. Replace with a cut down nylon bolt.
Remove remnants of broken peg mounting and tap hole in base
plate metal to take the nylon bolt.For indexing the mode switch:at
base of helter skelter cam the holes in cogs match holes in the
Alba DVD40 , 2003
(externally looks like Bush DVD2029)
5V rail reading 2.7V, resoldered the 3.3V SM rectifier marked
F131 960763 seemed to power up ok
Servos worked to index then 2 attempts to focus with
laser switching on and off but only short bursts of 0.3V
on platen motor, no continuous run and play.
Sometimes not that, just going into standby.
Jumpered door switch.
Cleaned lens , reversed end on end the
pickup ribbon, no change.
Removed KHM 280 laser unit (4 pin laser).
Removed 2 screws and lacquer that holds the ribbon connector board to the
assembly to inspect optical internals, nothing amiss.
Pulled leads through the ferrite rings to give more wriggle room.
2 adjustmment pots , nearest laser 840R left as set and other
was 436R reduced it to 305R and the unit worked fine for CD and DVD,
probably all that was needed perhaps only 400R or so required.
Alba DVD 40
Continuous fast flashing stanby and play LEDs and no focusing or indexing
if manually moved sled
disconnected mechanical deck ribbons and slight ticking ps
o.p ribbon read 2.8V,0,2.8V,0,10.3,0,-12.5
disconnected con2 no ticking
and read 5.2,0,5.2,0,10.2,0,-12.7V
Switch off mains and "-12V" stays much longer than other rails
no ticking top board disconnected
MT1379 34deg C over ambient
MT1336 22 over ambient
Awkward small DVD player with short ribbon. Requiring holding case open
vertically and trying to replace the wedge through the available access
space. Drops into the works somewhere and will not come out with turning and
shaking. In the process I assume its dropped on to the floor or something.
Has to be considered lost or very time consuming dismantling of everything
inside , in controlled conditions, and lost on floor anyway probably. So set
about thinking of how to make a replacement. Came back to the player for a
last attempt and the clip dropped onto the now under-placed sheet of
Anyway for next time , in earnest, any ideas on making a get-by
This one ribbon 13.5mm wide, wedge part of clip 13.3mm wide , wedge .3mm
down to .2mm over about 1.5mm.
Assuming near impossibility of making such a small plastic wedge with
uniformity across width.
1/ rubber, conventional or silicone, strip somewhere between .3 to .5mm
thick pushed into the slot with jewellers flat screwdriver. Then as cannot
think of source of small square section plastic, a matchstick laid over the
top and glued on the end cheeks?
2/ Its quite easy to make hotmelt string (when very hot melt) order of .2 to
.5mm diameter of reasonable uniformity over 20mm runs, just not
predeterminable , so select afterwards. Squashed into the slot but would it
relieve over time and loose closure force. Then match stick over top
3/ any other ideas asuming any salvaged ones you have, will be wrong
Reminder to myself for the next such awkward situation.
Like when refixing circlips I tie a piece of thread around the clip and tie
down the thread before squashing over the pin.
Glue a piece of thread to the ZIF wedge clip
The closure probably should hang in the ends but does not (broken) and even then
may grab onto the cable on inserting , so cable fails to bottom before wedging takes place
I did try with the closer wedge in place, located on end
cheeks, so perhaps nibs not broken.
It was not possible to put the ribbon back in through the gap without
pushing the closer in place before the ribbon had bottomed in the connector.
Begs the question - what is the correct method for removing these wedges ?
In a controlled manner so only 2mm of movement either end, and no more, so
the sub-mm pauls don't shear off in the process of overcoming the grab
A thought for re-assembly next time. A weak piece of rubber band from a
party balloon about 5mm wide. That would go around the wedge clip and hold
to the blue (usually) piece of celluloid stiffener positioned and held above
the lands section of the ribbon. When the ribbon is in place in the socket
then slide the closer down the 4 or 5 mm into the socket
3.3V on h/s of regulator, all connected and motors working drive but not optical ribbon, 12 deg C over ambient
on motor driver in play, representative DC of output leads
2.2V on L36
1.7V on L5.
LV on ps -22,-22,-11.5,.6,.5
Sometimes no LEDs at all due to bad 2 pin connector mains contact to ps
This DVD unit just used for the optical o/p checking of amps so not particularly concerned. It would eventually power up
with repeated on/off. ESR of all 40 or so electros seemed ok. Took DC readings of all
the exposed electro can ends and there positions and made a record for some other time incase
it showed up anything
Amstrad UF30 vcr
Failure to accept cassette
The sliding mechanism to lift the flap of the cassette was
floating loose so the lid was not lifting and fouling on the
moveable tape guides.It looks as though originally there
was a torsion spring and locating pin to return the flap-
lift mechanism that had broken and disappeared.
A linear spring seemed a better idea but no room for
a tension spring so fixed a light compression spring between the
sliding mechanism and the fixed part of the cassette handling
housing.Not very elegant as the spring bowed rather than compressed
but it did the job.
Amstrad VCR 4500
Failure to keep tuned channel due to worn /corroded
tuner switch bank contacts. If worth the effort
recondition as per tips file "ganged multiway push switches"
Amstrad VCR 4600
Jammed tape and no transport functions
Wide drive band from the motor drive of the mode
change mechanism had slipped off -excessive
axial play on the spindle of the worm drive cog.
Bent inwards the housing to reduce play,cleaned
belt and baluster pulley on motor and also scored
axial marks with the edge of a file to increase traction
on this pulley.
Amstrad VCR6100 vcr
Coin jammed in cassette carrier mechanism
All sorts of mechanical damage from owners efforts
to extract cassette.Cam train out of sequence and
tape-in positional latitude co-axial split pulleys
either side of carrier had broken retaining plastic lugs,
repair using small nylon spacers split axially and hot-melt
string glued onto plastic cog to allow rotation of the
captive metal disk/arm.
The tape-out/ingress latitude is provided by single
dual pulley adjascent to the worm drive.
Note the 3 limit finger contacts,tape in,tape out and
initiate tape ingress are not all activated in the same
sense from the cam followers.
Baird 8030 vcr
Playing too fast with "Pinky and Perky" sound
The pinch wheel was not engaging. The jockey wheel take off from the
cassette take up spool was jammed into the sliding arm.A circlip holding
this assembly had dropped off allowing the jockey arm to rise and then
settle back in wrong position.
Baird 8947 vcr
Cracked pulley tyre on supply reel satisfactorily
replaced with circular section 7/8 inch "O" ring.
But main problem not found. Sorting out the transport
and checking there was no video or sound output
only modulated carrier (test signal). Large
crack in main pcb in area around the central plastic
hold down anchor. Repairing them the vcr entered
an illegal mode with capstan running in reverse
and would not power-up from standby. Never found
the fault and abandoned.
ps voltages when removed from vcr
?,?,0,0,?,?, 12,21,.7 There was voltages on
pins 1,2 ,4 and 6 when connected probably gated on by
control voltage on pin 9
BUSH VCR168 = GOODMANS GVR3400 VCR
Tape jam with tape wrapped around spindle and pinch wheel and large
loop of tape in body of m/c.
In play and FF one of the cogs on the jockey/idler assembly rubbing against
chassis,consequently removing drive from the take-up spool, so removed
jockey assembly and packed out with a thin washer.
Baird VC121 vcr
No REW or FF.
The set/reset mechanism on the main slide bar section of the
mode change mechanism was not going to set condition in retracted
state due to wear on the plastic pivot so shaved off a small slice
of the moving dog that engages with the slide.Thought the spring action
was a bit weak so replaced with a beefier torsion spring but this
led to the mode switch motor stalling and when going into STOP
mode would unlace tape then go ingo REW.So settled with original
spring.After removing the cassette carrier mechanism temporarily
fix the main deck to the chassis before turning upside down the
Canon E90 camcorder
C2901 and C2931 panasonic SM electrolytics had failed and
electrolyte had damaged some tracks.Replaced the 47uF,16V and 100uF,6V
with conventional caps. Remade trace between pin2 of IC c555 back
to CN29093 connector and replaced IC protector RR2933.On reassembly the
remaining 3way connectors are line plug and socket to the mike.When
replacing the cover beware of disrupting the 1mm spacing ribbon
cable that lies under the top cover (cover with the R/C window).
Locate this top cover on the deck side side cover and then
slide the viewfinder cover onto this top cover.
If this is 1990 video8
To open cassette compartment, motor is "above" top left of video drum , 5V
colour reversed polarity will open the carrier (needed a video8 tape)
Daewoo GB14F8T2, VCR + TV, 1999
Chews tapes because suspension arc for the jockey-wheel/idler/ slip
clutch assembly is an arc of plastic moulded over the steel of the chassis.
Then has cracked and catches the carrier as it swings across . A web of
plastic filaments joining odd bits of plastic pillars etc, located by
flowing through holes in the metal at formation, but one of the
interconnects forming this arc. If anyone remembers all those Philips car
radios that used plastic rim to the metal capstans, that failed, because of
excess heat inside the dashboard of a car in summer and differential
expansion of incompatible materials. Or I thought so ,this in a domestic
TV/VCR combo with ps and heat over the VCR , not under. This is a Daewoo
GB14F8T2 using coincidentally? lots of Philips chippery. So could it be just
an age effect, 1999 vcr, this hard plastic contracting with age
47R ,10R (in cct) ,7808 vcr section G3 on the deck
Remove screws holding vcr base plastic tray to the
steel over cover. With TV face on protected surface , remove the vcr section and
tie the tv steel work to the outer casing to prevent it falling against the
CRT neck. 2x 5 way cables P405 to PA405
2x 3 way cables , speaker to socket near headphone socket.
Both 5 way not needed for vcr use, but 3 way cable is required.
Slow tape rew , light getting to takeup side end sensor.
Broken click switch , one of pair, robbed standard click switch
button and dome and hotmelt to rear rather than 4 broken pips.
TV 270,1R,1K,3x8K1, 5k1,10K,
2SD2499, 2x 5k6, 1,47,r33,47r,2x 220k,3r3 10w,100k,8r2,27k
TDA8351, 7812,78?? + another, TOP210PFI
7806, S5707, 3x TDA61060
Whine and occassional "emergency" on screen , requiring switching off
and eject /entry cycle.
Take up spool brake broken in that felt had dropped off and migrated
under. Cut some felt and hot melt on.
To clean supply back torque band , slip over the spool and twist
swivel anchor 1/3 turn
Slippy slip clutch
To separate deck , 5 screws. Remove jockey assembly from top.
Mark underneath slip clutch assembly parts before dismantling.
Screwdriver blade to separate central parts
Added 2 rings of strimmer plastic , 25 mm long under the central
heavy spring, 2mm spacer on closing. Reinforce drum ribbon
end before reinserting. Emergeny statement, latch up but not tape mangling, maybe due to failure of
pinch wheel contact. Cleaned and pushed a piece of strimmer plastic
between anchor hook and tension spring loop to tension a bit more.
Seemed to occur at beginning of tape, switching off and on seemd
to be a getby cure to that problem. Maybe structural twist thing as never occurs
when connected up outside its case, only when VCR and TV and steelwork
screwed together .
Cassette in eject position while lifting flapand reintroducing
assembly into the front of the case.
de Graaf / Granada V/WHSJS1/A vcr
Transport whine and occasional failure to play
Excessive slip on slip clutch and worn capstan bearing.Cleaned spindle
and introduced graphite powder as temporary fix for whine.
Ferguson 3V57 vcr
Nasty noise from inside.
The cassette loading motor running continuously. The triple
cog assembly had failed due to one of the nylon locators
shearing off and another half broken.The half broken screwed
up from rear with small self tap screw and epoxied in place.
Repeated same with 2 still good pegs. The failed peg used a
bolt epoxied in place with head ground down partially so
as not to interfere with the intermediate cog. There are
index dimples on the cogs to align on reassembly.
Ferguson FV21R = JVC HRD400
Poor FF and then tape jam.
Distorted spring and graphite laden disc inside
the idler assembly ,while awaiting a replacement idler
tried reassembling with the spring replaced with a disc
of squashable rubber 0.2mm thick and all transport
functions worked ,whether long-term solution,not known.
Intermittent tape transport failures.
The idler assembly was riding high and interfering with the surround to the
drive motor spindle. Also the aluminium hoop that retains the spring
that holds the idler in contact with the motor spindle was just high enough
to sometimes touch the bottom of a cassette
Manual record and short lead-time timed record OK but failure
of overnight record.
Replaced the the 3 small electrolytics on the clock / timer panel
and broke / inspected /remade all connections and
also at someone's suggestion tightened all earthing point through-pcb
screws on the various boards.Somewhere along the line must have
effected a cure.
Ferguson FV21= JVC HRD170 to 330 vcr
Jammed mode switch actuator arm.
The jockey-wheel assembly was jammed against the 2 steel
pins of the actuator arm,access via removing the two motor mounting
screws.For alignment of the mode switch the 2 V-shaped notches in
the body of the switch are aligned when the m/c is in neutral
with no motors active.
Ferguson FV21R VCR
A few mis-tracking lines at bottom of image
Dirty and compressed felt on the tension arm
copper/felt belt wrapped around supply reel.
Mark the position of the fixed anchor and then undo to release
the band.Clean the felt with methylated spirits and slightly
roughen the felt and replace.More permanent replace the felt
(from haberdashers).Pare off the existing felt and glue (as in
hints/tips gluing plastic) a new strip and razor back the excess felt to the
copper band..Refit and adjust for correct back tension.
Fisher FVH-P710 vcr
Poor FF and REW,squeel on mode change,eventually jammed tape.
Stretched band leading to FF and REW problem.Stretched band on mode
switch assembly leading to remaining problem.
Fisher FVH P906
No tape functions.
Original fault was excessive slip in the slip clutch under the supply
spool but someone had tinkered in area of 2 locus cams.There was greased
track in the smaller of these cams but no locus follower assembly.
There is a long pivoted arm that engages with the large locus cam but needs
another linkage to activate in FF and REW.There is a return spring to
this arm but introducing another stronger spring pulling in opposite sense cured this
problem.What the original mechanism was I have no idea.
GEC V4006 VCR (1986)
Tape stuck in machine.
Goes into safe mode if timed-out the mode change
from tape in to tape play. Due to gooey grease on
the worm gear of the mode change motor.
Spun up with brush loaded with meths to
clear away grease,then silicone greased. I think
the motor could do with more voltage in this last
sequence - it should not need grease on this cog IMO.
Loose wires in mains plug had led to 400mA internal
mains fuse blowing
Goodmans TX1200 vcr
Low sound for self record ,normal level replay from pre-recorded tape.
Traced the audio component feed back from the audio head,as distinct from the bias oscillator feed
which was fine 30V or so. In the audio trace near the bias adj. and osc. transformer
there is a 22K sm in line so paralled a 13K sm with this to raise the level on record
Goodmans VCR 2000
No mechanical functions.
The main deck spindle driving the jockey wheel assembly had deformed
sufficiently to break contact with the driven pulley bottomed out
against the arc slideway in the chassis. To get to this larger pulley
remove the take-up spool and disengage the torsion spring retaining it
under the chassis. To access the small pulley - remove the belt drive pulley
under the deck. Remove the cassette carrier assembly and 2 screws holding
the spidle housing for this drive assembly. Could not find a replacement
barrel type pulley tyre so removed and packed out to take up slack - a
bit lumpy-bumpy but hopefully would bed down rather than get more noisy in
Godmans VN9500S, 2000
No functions, showing Err on display
On loading the mode select gearing does not stop at selected mode
and continues to the end stop.
Dismantled the mode switch and reconditioned as per pots.
Index marks on green and blue sections of mode switch
are adjascent for tape out position.
Mode switch connections at connector
Tape out , D-Gnd
Full FF, not half cock, B-Gnd and B-E
Play , E-Gnd
Granada YH2 vcr
Only forward,no rewind or play,no response at all eg motors etc.
Dry joint on the signal conditioning circuit for the supply-side
end of tape sensor.
Grundig VS 520GB vcr
Jammed cassette carrier
This is the same tape deck as used on many Panasonic models
with the helter-skelter cam for moving the pinch wheel. On the
main drive side of the cassette carrier the two part lever
mechanism (that through connects to similar mechanism on left side)
had bent so allowing the driven pin to disengage.
Beefed-up the lower of the two levers with a piece of metal strip
hot-melt glued in.Clamp the lever of interest against the body of
the cassette sub-housing.Gradually (to avoid any more distortion)
melt the glue into the recesses and pre-heat the metal strip to bed
into this now cooled glue.
8 or B80620 on drum
VHS 4 head system
To remove drum
Desolder non obvious set of 4 pins laying circumferentially, not radially,
slide out of holes in the black. The black plastic must move with the drum as you lift it.
If it jams cut the other pair of terminals so as not to lift the mini pcb of the brass of a head
and break the fine wire
Type name of 2/3 size VHS video head drum?
Instead of 62.1mm diameter, 41.3mm, as used in standard size VHS cassette use camcorder Hitachi VM 500E of a couple of decades ago.
Used full width tape, not 2/3 width ie 8mm, AFAIK bulk standard VHS tapes
maybe fairly straightforward compatibility. there are 4 equally disposed heads, on 2/3 diameter , maybe run it at 4/3 satandard drum rpm and 270 degree wrap
Anyone know where to find the dimensional specs/standards of helical scan systems (before wear anyway) particularly diameter between head face surfaces and angular offsets , particularly for DAT and video8 systems?
By not work, pre-recorded Bond movie on a DAT tape plays a useable B&W image but absolutely no sound . In 6 chanel PCM audio mode no sound on any of the 6 chanels of a tape with n6ch audio recorded to it
Grundig VS520 vcr
Lock-out from the system-no tape functions
To unlock the code-lock: Open the vcr and find the testpoints on the display`s
pcb. This is usually 2 pins near the upper edge near the keypad,near the 0 keypad.
Connect the vcr to the power-line, short the test-pins together and press `stop`.
Setting the head-switching pulse: Insert a real prerecorded cassette and
short the testpins. Press `6` on
the keyboard. The vcr will now align the switching-pulse. If successful `A6` in
display disappears and
0 appears. To exit the service-program, press eject or stop.
To reset a timer alternately press the timer button that flashes and `0` until it stops
Hitachi VT 35E 1984
No tuner and poor FWD & REW
Hank of wires going to tuner were draped over the top edge of power
supply board. With owner putting stuff on top of machine meant one
and half wires had been severed. Transport functions returned
Hitachi VT 64E
Erratic tape speed.
Worn out capstan motor,swapped with the similar looking
motor from an Akai VS4 but ran too slow presumably by
a factor of 7/12 ,had to swap the 48 pole tacho
with the original 28 pole tacho.The tacho housing
is just pip-swaged / welted to the body so easy to
Hitachi VT220E vcr
Gradual fading of colour and intermittant loss of colour
Dry joint on the SAW filter also bad solder joint of a SMD
chip mounted on the stand-off sub-board located near this
SAW filter,unknown identity as coated in black epoxy paint.
Also for good measure the tape transport would have
soon failed due to compressed / perished pinch-wheel.
Hitachi VT 430E
Intermittent failure to play, just a judder from the motor then drop out
to safe mode, then later intermittent colour problems eventually intermittent
total loss of colour.
Drive problem associated with black epoxy painted daughter
board on main board near SAW filter.
Colour problem associated with the other such ceramic sub-board
near SAW filter and under tuner sub-board.
Pressing on the main SM chips on each of these boards 'cured'
the problem. Not knowing how to remove epoxy paint , stuck
a rubber pad to IC on one and top of crystal on the other and
tightened a couple of small cable ties around and under
each of these ceramic boards.
JVC DR MH20SEK HDD DVD PVR, recorder, 2005
Ticking smps , flickering HDD and DVD lamps , no display, no fan, no standby, no functions
Add 470K over the mains DC cap for personal protection if you
forget to discharge it.
Putting LED over the OC LED stops the SMPS driver.
Adding a 200R to this LED back to C5305 , would be on most of the time.
5.6V there and Vref of 2.4V on the TL431, 3.7V on TL431 out.
Added 10R between G5563 and transformer.
5.8V and 2.1V on either end of the introduced 200R to LED
13.5V over C5103 not enough
Ticking caused by failing hot side electro on the supply V for the
switching ic driver.
Uprated and smps is stable putting out right sorts of
higher DC , according to the schematics.
Good 5.8V on 5.8V rail
No 5V or 12V on HDD supply
12.2V and 5.8V on D5307
12.6V on C5304
-29V on R5397
-4V on fan
-11.5V on CP5301
6.5V on CP5302
6.5V on C5307
but 0V on D5310 and 0 on pins 5,6,16 of IC5302
But the LV 3.3V etc regulators seem to be waiting for a control signal.
Still no display, no fan , no button function and flow chart on e-service is in
Russian but I suspect like the error code page needs the display up and
Broken land number 23 on the main ribbon meant 17V supply was going nowhere,
not me honest, I always wrap proper cloth tape around these ribbon ends , to
beef them up, before inserting them. Someone else had tried repairing this
one. All functions seem to be returned .
SMD 565A, 665A and 563D, logo of a fractured italic M for Mitsumi
are MM1665AH, MM1565AF, MM1563DF and MM1663DH
Fitted a mains switch under the front panel on the chamfer , as "off" consumption
measured at 15W. Beware of trapping ribbon at DV-In , under the DVD
15W or so in "off" state of this unit but by going into settings and selecting
"deep sleep" mode then 5.6W in this mode with the clock display off, 24mA
mains current and by rotary meter
JVC RM-SDR008E most functions of RM-SDR011E but not power/standby
or delete function
JVC HR A630 VCR , 1996
No display, no finctions. HV DC at the ps but no oscillation, 5V
at the HV side of the opto-isolator.
I'd removed both the switching power FET and the biasing driver 2SC3616
because there was a 220K resistor between G and S which was not on the
schematic of JVC HR J725 with similar ps I was looking at and confusing
the cold checking.
I removed the 2uF,50V to the G as well, nothing seemed to be wrong with it
but was intending to replace with 2x 1uF,63V in parallel but accidently put
them in series so only 0.5uF but it worked all the same. I'm not convinced
it was a problem with that electrolytic but its the only thing I actually
Its RTFM for setting the tuner, also for HR A631
Ch Set/OK/Jog Up/down / Ch Set
For manually resetting
Ch Set/Jog /OK/(number)/OK or Store/repeat/Ch Set
Fine Tune, select ch/OK 3 times/Jog/OK
"Guid" relates to Videoplus
JVC HR A630 VCR
re-use after storage, immediately blew internal mains fuse.
Blown smps 2632 s/c all round,replaced with a IRF740, with insulator and
one of the locator nibs ground off the heatxink plate,
but running hot, so obtained a 2SK2632 replacement .
JVC HR D170 vcr
Poor picture in E-E and play and "test" card
The owner had been using with the TV tuned into a strong
secondary image frequency of the output modulator.In this
m/c there is a 2-sided 3-hole corner brace over the p.s area
which was loose (no designed fixings) and nearly touching
an interboard connector point on the top pcb,also could
have slipped into the p.s. area so bonded this corner brace
to the 3 plastic support lugs.
JVC HRD230 vcr
Intermittant failure to record timed recordings
Replaced the small drive belt that powers the
mode change mechanism.
JVC HR D400 vcr
Warbley sound and switching line in image.
The switching point could not be shifted off screen
by adjusting R430.Fault was poor contact/s? in the
connector CN1 taking drive signals to the deck motors.
JVC HR D400 vcr
Mode switch alignment.
This switch has 3 pins ,the only position where all the pins are s/c
is in the neutral position at the extreme position with the slide bar nearest
the locus cam.There are numerous positions where the 2 outer pins only
are s/c.On the locus cam the neutral (index) position is when the innermost
small hole in the cam aligns with a similar hole in the metal cover.
No play mode
Capstan wheel not rotating.Due to failed connection in the
nasty 1.5mm spacing ribbon connector socket on the mechanical
deck.Relieving spring action to release cable ,resetting pin,reintroduce
cable and have disturbed another line.Redo and this time break one of the
leads.Sod this for a lark.Desoldered the socket after cutting away some of the
plastic of the chasis.Soldered the ribbon to the top side of this socket,beefed up
with hot melt glue and soldered back in place.
JVC HR D560 vcr
Both video heads were dirty so cleaned (power off)
with stamp-sized piece of clean copier paper.
For thicker accretions dampen paper with methylated spirit.The trouble
was probably due to the flimsey foam cylinder on the
auto cleaning arm breaking up.Removed the whole cleaning
arm assembly ,never have liked the snatch loading in the
most sensitive area of a vcr when these arms engage with the
still spinning head drum.
JVC HRD610EK vcr
No control over tape functions,only choice of (play+FF)
Lack of freedom for click switches behind front panel
meant FF and REW click switches one was deformed
and the other was intermittently on.Replaced these two
click switches and packed out between pcb and stand-offs
with 2mm spacers.
JVC HR J225 , 1995
Mechanical deck much as HR J725 or HR J825 so applicable there.
Mangled tape due to eratic forward and reverse tape wind.
The cogs not meshing due to wear inside the complex
assembly over the main large drive cog.
No need to remove deck in retrospect.
But for replacing main drive band etc
Undo the 4 ribbons and reinforce the thin one before remaking
2 large screws and 2 nylon pcb standoffs at front.
Remove metal band along rear edge of rear slide bar.
Remove the metal strip between the 2 intermediary double
slip clutch cog drive spindles.
Remove the circlip at the wormdrive cog , but mark
the position , in neutral, just in case and bend the
white slide bar up and rearwards to get at the central
pivot. Beware the spring buried inside, only 50gm to
compress to half length. Replaced it with a stiffer
spring to keep the radial arm jockey assembly from waggling
up and down. Required reinforcing the slideway retaining
claw by gluing to the central IR light guide and
cutting back the corresponding hole in the base
of the cassette carrier to stop it fouling on the light guide
as it bends a bit.
The end of tape sensors also use light guides and the
tape-in switches are below the CC display.
Drop the cog on the end of the swing arm by padding out
with a pair of running washers.
If the top plate , between intermediary slip clutches,
fails to locate properly , check whether you've
doislodged the fiddly black swinging ratched that engaged with
the supply spool. Unplug the ribbon to the function select
motor before powering with 9V battery , in either direction,
to check for any jams/ slippages with both slidebars etc.
The cassette carrier is easy to remove/reassemble if
in neutral position and tape out, just 4 screws, no
wires or index marks to observe.
JVC HR S5000
Jammed tape with small motor spin noise.
Stretched mode switch belt.
With m/c powered on manually assist the driven side of the
mode switch until the cassette ejects.Check there is no movement
of mode switch between powering on and off on the front switch.Mark the
positions of the cogs on the mode switch gear train.
To replace belt.Desolder the ribbon cable,remove mode sw cover screws,pull
away leaving the through chassis linkage in place.Remove plastic
circlip on end of spiral cog and bend out end housing.Change belt and
reposition train to marked position.Reassemble the mode switch while pushing
the rack part of the mechanism to position in direction of mode sw.
P.S. voltages in stop mode on 9 pin conector.
1 -36V,2 54V,3 0,4 7.8V,5,6 0,7 17.7V,8 0,9 15.7V
JVC Nicam stereo (model not noted)
Non ingress / egress of tape because corner
of vcr housing is damaged by a knock.
This repair is probably valid to most vcrs with
plastic actuator triggering the cassette entry flap.
The knock had broken off the small spigot that touches
the flap.Cut off a small piece of right angled aluminium
heatsink and abraid off the surface to be bonded
to the remaining part of the plastic actuator.Hold with
pliers and heat with a hot air gun to melt hot-melt glue
to give a good bond to metal.Pretreat the plastic surface
to be bonded (see hot-melt string in tips).When ready to
bond then heat with soldering iron both parts and hold
together till bonded then reinforce bit by bit around
LG LH-C6235? , 2005
VCR + DVD player
will play DVD only if assisted start with a finger.
Remove 2 sc to remove top clamp and cleaned lens with meths
Matsui VP 9031 vcr
Lack of colour on off-air signals
Something wrong with the tuner ,owner accepted "cure" of plugging in a UHF
Fails to play
Too slippy slip-clutch.
Remove plastic circlip from underside, assembly is
sprung. Separate both halves by placing a blade
in central section containing 2 concentric springs.
Clean and slightly roughen.
No functions followed by shut down and no
tape eject. Loop of tape in m/c.
Seemed to be problem with mode switch or
tape out sw but not so ( action is through deck and push sw on the
main pcb. Capstan spindle rotates
but not the slip clutch. On removing the base cover
a small piece of black plastic was loose and the drive belt.
Plastic split radially and dropped off the end of the capstan
spindle. Glued back in place ignoring the split.
Switches off after a few seconds and no functions
with tape stuck inside.
Push the sliding part rearwards on the carriage drive rod
and turn the bit with the 2 fins to manually lift out
A bit of gummed label stuck to the hole uinder the
cassette interferring with the IR LED post.
Poor jerky play/record towards the end of the tape
furthest from the start end, also poor wind
when most back torque on the tape spools.
Lightly oiled on the capstan spindle bearing.
With test tape in and exposed tape spool.
20 to 30 gm at 38mm radius hole on the spool
corresponding to 18 to 22 gm at the slip clutch.
Upped to 50gm at radius 45mm on the slip clutch
pulley by packing out with matchbox and rubber
band and turning by hand. Gave too much torque to
measure on the test cassette spool.
Changing the main drive band to a larger
cross section and tighter one led to failure
of play function due to misaligned capstan spindle.
Nasty noise , no functions and switches off after 5 seconds or so.
Broken mode switch motor band.
On replacing the belt, B&W picture and multiple
Due to something mispositioned after trying to
manually drive the gear system to eject the tape
stuck inside when the band broke.
Matsui VX1100 vcr
Intermittantly fails to play or REC.
Could not induce the problem myself.
The slip clutch assembly had obviously been rubbing an
arc on the take-up side of the deck. Removed the slip-clutch
arm and packed underneath with a thin plastic disk. Cut back
the small proud ring on the top of the arm to relocate the plastic
Matsui VX820 vcr
FF and REW but no play also intermittent non eject
Once again that cross-linking pin that moves the take-up
tape guide into play position had dropped into bottom of
machine due to cracked plastic so tape guide contacts
the pinchwheel.Replace the pin with a longish(for a 8BA bolt)
the correct length and beef up the plastic with hot-melt glue
and cut back a bit of the take-up spool brake so it does
not foul on the beefed-up linkage.The intermittent failure to
eject was due to the owner placing TV immediately on the
vcr and deflection via the small rubber block on the right
hand side of cassette handling mechanism and distorting
the housing.Also rattly squeal from the anti-static
conductor rubbing on the head rotor spindle was cured
(short-term?) by a very small dab of silicon grease.
Matsui VX880 vcr
Loud rattling noise in FF and REW
From underside remove the complex slip clutch assembley
that drives the jockey-wheel .Pack out on the
end next to the plastic ciclip with a small graphite loaded
plastic running washer to take up the wear.
Matsui VX880 vcr
Intermittant failure during play
Cracked pinch wheel surface
Matsui VX1108, 2000
Jammed tape, for no known reason
The nearby phototransistor survived me putting
9V across it both ways, thinking I was feeding the
load/unload motor, thinking the wires went to the motor.
Removed and reseated leads and plugs/sockets-under and no further problem , so
perhaps just that.
ps 82K,220,1R8 and LV side 22,220,1R
Osaki VCR 34H
Intermittant no REW or FF
Originally thought to be slip clutch problem.
To get to remove cassette housing and rubber drive band
underneath - to assist reassembly cut away the plastic chassis
around the bottom pulley driving cassette carrier.
Remove large bottom pulley from the slip clutch assembly.
Long armed intermediary double pulley assembly is on
the right and is engaged in neutral, both arcing leftwards
for REW or FF. Small sliding bar interfering with the right
hand brake should be fully retracted in neutral.
Double slip clutch assembly order of assembly
Small black cog, spring, steel disc,white cog,
large black cog.
End of left steel slide bar goes under the end of the
long armed arcing dual pulley assembley.
Trouble became apparent only when operating with
the cassette carrier removed but ribbon cable connected.
The doubled up sliding white slide bars have a trigger on
right end that is delatched when kicked by the "star" cog spinning up.
On top of this double white slide bar is a latch that sometimes
stopped angled rather than inline and needed the return torsion spring
increasing the torsion.
Osaki VCR 34H
bad jerky play / rec
Slip clutch felt replacement
Well I've always called it felt but this is worn down to a woven backing
that is well glued to the pulley. I suspect it is more pool/snooker table
baize material rather than haberdashers felt, but why always white ? some
special formulation. Unobtainium full clutch assembly so renovation of what
Anyone know how to make neat punch-cut dual concentric cut rings without
proper setup. Small central hole ok with normal leather punch ,but outer
Probably I'll cut the centre hole in some green baize, glue to the remnant
backing (or start again stripping of the old ?) , compress as its curing and
then cut away outer excess with a scalpel.
Have decided to punch the small centre holes, then glue onto the original
surface as no integrity left in the woven remnant matrix. Then an adjustible
craft cutting compass , as parallel axis of cutter and centre it will locate
easily here, to cut the outer ring.
These compasses eg
are great for partially cutting rubber sheet to drive bands. Partial, so to
keep the cuts regular and concentric and so build up a large range of bands
from one sheet, then finish off the cuts by hand with scalpel , guided by
the initial cut.
Cutting compass worked well, almost looks a pro job except for the green
colour. I suppose only one world supplier of whatever this slip clutch
material is. Sorry. I don't know where you get baize from, my source was a
pool table repair man working at a pub and was not averse to a drink for the
old cloth in his van.
Worn out both slip clutches in jockey assembly
To undo, plastic circlip to remove drive pulley. Topside introduce a
dummy cassette to release the front 2 sc then the internal 2 sc
Loose yoke sarm in the space left side top over the left jockey wheel needs to
go under the arm.
Without large drive pulley in place ,add middle sc to top clamp.
Introduce dummy tape to clear the front 2 sc holes to replace them.
now works in play, tape-in , tape eject, not FF or REW
sc from slideway on the under surface protruding and touching a cog, so
packed out the sc with a washer. Probably wear allowing to move and wouild
eventually touch the chassis.
Left swing arm jockey assembly has the smaller secondary cog
Bad play/rec tracking due to grimey supply side band brake, cleaned and
adjusted the setting
Osaki VCR 34H
Fails to eject etc, dropping into some undefined state.
Repeated pressing power on/off would eventually
cause a click of a latch and then normal functions.
Failing belt to the second largest of the 4
Slipping at the pulley that strangely crudely engages with the rubber band around the
capstan motor pulley. This action disengages the compound slide mechanical latch
on the other side of the deck.
Osaki VCR 34H
Replaced pinch wheel with one 1mm greater diameter,
as slipping, tape damage.
Osaki VCR 34H
VCR works fine in play and record as long as the head drum gets up to speed
by flicking by finger while the tape is lacing.
So assuming nothing wrong with the bearing , normal free movement when
Sometimes drops out in play or record on hitting ruffled tape so a motor
3 lines from the TA2620D driver IC to the coils and further ends of each
are joined together but not returned to anywhere so presumably 3 phase.
There are 12 coils from these 3 lines mounted on the iron frame with 12
The surrounding magnet rotor part has 8 North poles, 8 South poles.
On pcb 3 hall effect sensors presumably for relative phase monitoring these
magnets and a separate magnet pip on the outside of
the rotor and 1 hall effect sensor for 1 pulse per rev.
But what is the PCB track/coil ? etched on the board under the rotor.
Not connected to anything else on the motor sub-board and just copper track
as far as I can see,
roughly like Beta vcr tape path, one track runs right round circularly and
returns 'circularly' with 24 castellations pattern.
What and how is it sensing or sending anything? and what's its function ?
I tried another matching salvaged multimagnet rotor and can eliminate
magnetisation loss as reason.
The castellated track seems to supply a tacho signal. The off "main beam"
multipole magnets on rotating must be flux cut by the castellations inducing
a low level signal. May feed a drum loading/cut out as well/instead of
monitoring drum motor current - i don't know, not followed path.
Main problem seemed to be ps problem. Caps ok.
Decided to uprate .1A biasing Tr (2SC1740) to .2A BC449 swapped legs, and pass
Tr from 3A to 6A TIP41C and
changing 13.6V setting zener (marked C12) to 15V to see what happened. Much better start
up torque, checking Vs I'd put a 13V Zener in by mistake
and drive volts was less at 11.5V down from 12V. Replaced with intended 15 V
and drive V then 13.6V but very little torque.
Returned to 13V zener and no problems since, caters with rinkled tape now
without cutting out, i doubt its a permanent cure though - weird.
The zener on the drum driver board seems to be 2.1V, i changed it to 2.5V
and the whole drum drive ceased, changing it tp 1.5V zener the drum speed up
was good and cync-lock was good but stayed on permanently and left it
in this situation. VR132 changes rotational cync point of drum .
Cutting the line to the 24 "castellation" track meant no cync-lock with the drum running up
to maximum revs, so speed feedback to drum servo cct.
Excluding microswitch lines the lines to this drum driver board are
(before altering so maybe wrong)
To remove the 3 sc for the slip clutch introduce dummy cassette or trigger
both leading catches of the carrier , when powered up, to clear the
front pair of sc and eject to allow removal of the plate.
Well I've always called it felt but this is worn down to a woven backing
that is well glued to the pulley. I suspect it is more pool/snooker table
baize material rather than haberdashers felt, but why always white ? some
Have decided to punch the small centre holes, then glue onto the original
surface as no integrity left in the woven remnant matrix. Then an adjustible
craft cutting compass , as parallel axis of cutter and centre it will locate
easily here, to cut the outer ring.
These compasses eg
are great for partially cutting rubber sheet to drive bands. Partial, so to
keep the cuts regular and concentric and so build up a large range of bands
from one sheet, then finish off the cuts by hand with scalpel , guided by
the initial cut.
Cutting compass worked well, almost looks a pro job except for the green
colour. I suppose only one world supplier of whatever this slip clutch
material is. Sorry. I don't know where you get baize from, my source was a
pool table repair man working at a pub and was not averse to a drink for the
old cloth in his van.
But poor tracking was due to too much supply spool braking. Dribbled lubricant
down the spindle of the spool. Added a loop of wire to the tension spring of the
swing arm to reduce the force, cleaned inside brake felt surface and all other tape
engagement surfaces including drums.
Sometimes REV/FF would not operate, due to protrusion of a slide way mounting sc
end , touching one of the cogs. Presumably wear alowed the cog to drop, padded out the
sc with a washer but eventually the cog will rub on the deck steel.
Of the 2 cog bearing arms that drop away on removing the cover plate the one
with the smaller driven cog is the left one. The loose yoke steel goes under the arm of one
of these cog assemblies. If this arm gets out of place then tapes with the write-protect
tab in place may not play or record, reposition or fit some thick tape over the tabs
Panasonic DMR EZ25
Long wait between returned display rreturns
motor/stepper jerks etc
ps voltages agree with manual
slow to respond to buttons like open/close and will not play discs , will go
through the sledge/platter sequence before ejecting, laser light present.
Most of the time just says please wait.
Will not respond to the almost inhuman hand-span 3 button push at power-up
to go into service mode.
Will respond to forced power off of 10 second push of power/standby button
and go to presumably flashing clock of 0:00 that does not advance to 0:01
I've reseated all ribbons and ESR'ing caps is ok in main board control area and the top
Well that was an interesting exercise in fingertip manipulation . Fathomed
(accidently ;-)) how to release the central area of the deck then remove
the 2 sled slide rods and then remove the laser assembly. Now know it is
type RAF3332A. Then get inside the optics, all without disturbing any
set-screws for slide or optics I think, and reassemble. The 45 degree mirror
was filthy with dust and even a curl of hair right at the centre, if it was
the outlet lens that was that dirty then a CD/DVD player would work. Not yet
fully reassembled and powered up yet
Just for future reference , not blown , but anyone know the full series /
letter rating , assuming some sort of standard
1.5mm long SMD, with letter G , measures 1.6R ? IP7501
and one marked P is 1.2R ? ,IP1501
the letters P and G are marked in large letters next to the otherwise
unlabelled devices on the overlay.
As in Panasonic kit , perhaps these 1x .5mm SMD , with solder over probably
looks more like 1.5mm, The ones before me the ceramic core is light blue
colour, maybe relevant
Rated Current (A), Marking code
light blue body with white letters and a printed brassy line between pads so probably not those pana PDF fuses
larger SMD IP on the digial board , black letters on white, this time N and K
Fan never runs , on-demand? 12V supply to its contro area ok
To release ps , switcher MR4012, unclip main ribbon and pull up, 2 sc at rear, 3 pcb sc
Dc/DC SMD converter marked 1014 659
othe larger SMD marked 1717?
With DVD deck top cover removed the 4 noise sequence is the sled moving , single noiuse is with
platter kick also. Laser comes on and focus servo works.
Cover has part of the mechanism , a slideway on it. If in wrong posistion remove , let
mechanism settle in no-play posistion and return the cover.
Fan ok on bench ps
ST 24C64 supply 3.3V
4.8V on supercap
W52 and W53 at main interconnect for I2C IC7404 , 24C08
In the deck 2 slotted optos and micro sw associated with main white rack/cam/slider bar
To change drive band , create a gap by pushing this slide bar further than normal by
bending a end-stop , then the rack will lift enough but don't remove.
The slotted opto levers need joggling to reseat the slide bar.
There seems more movement available at the other end of motion of this slde bar but something
stops it. That is another larger end stop , in the play posistion, near the motor.
Move that to allow the slide to move further and the dog in the slideway can move into the
extra track , for maintainence, and the whole central section will drop out , so be eary.
Fro the silver steel rods away from the platter, bend back the rearmost flaps and remove
both rods enough to slide off the laser unit. The spring does not dislodge from the helical
Matsushia logo and 3332A , 32F, 6831H, so RAF3332A
1 sc releases Al cover ,
under is 2 long sc that hold the lens and focus unit to the h/s / pcb
Beware the arm is sprung and just an epoxy glue spot, heat to break the join.
The other epoxy spot near the ribbon conn breaks from the pcb easily. These epoxy spots
seem to be just thread stop and not precise locators.
Cleaned lens with a small twist of cotton wool and meths. Careful of 6w ribbon
connecting the parts as left in place. The gry Al-looking bit over the internal lens is
probably grey sticky foam tape keeping it in place
28 pin IC marked
Forthe green ribbon connectors , with the clip open there is no green to be seen
through the gap , push the clip home and an even width of green should then be seen.
When powered up with SD and HDMI boards seemed bettewr behaved in that
a Disc will remain inside happitly with out ejecting but still U61 and no read of disc , also with presets
adjusted back to as found
In the optics th ealuminium grey looking "clamp" over the internal lens is sticky grey tape.
Desolder the 6 pads to the focus coils. Heat epoxy spot undr the presets board
before lifting away. 16 w foil ribbon to diode array .
Remove clip holding in the filter , gouged away glue holder th eblack ring with filte rinside , to remove
?, picked away at glue spot around the Al frame that holds the laser until it would lift off window clear
and interior grey clear not burnt, uniform grey, uniform sintered / crystaline
appearence tothe active face
For self-education I decided to take apart a presumably defunct DVD laser
unit to see the internal layout of splitters etc and what the laser diode
looks like. Googling in images I only find graphics or external views,
anyone know of a WWW view through the window of some good laser diodes ?.
Looking via x30 at what I assume to be the active surface is a uniform grey
colour and uniform crystalline appearance rectangular face (ie no obvious
hot spot/s) like a finer grained version of a fractured surface of cast
aluminium or monkey metal.
Is this the infamous Catastrophic Optical Damage ?
This one has a marking of RH on the body.
So if the active surface of this pic , (different internals to the one I
have looked at )
is the uppper part of the die then the same grey granular appearance
Smaller than the resolution of that photo or the small blackish rectangle on
the end of the "bar" that ends flush with the granular surface, on a line
between the centres of the end faces of the conductor pins.
Dimensions of that rectangle about one of those wires diameter x about 4
I went the "unsafe " route and put 2.5V and 560R over the 1.6V DVM diode
test junction ( 1.3V of the other pair of pins of this recordable DVD
laser). The dim and dull-red light came from the junction between the grainy
block surface and the pillar . In my one the end of the pillar is a
"polished" surface with no metalisation over it unlike the one in that pic.
Splitter cube between laser filter and foam cover s lens. Either side is a cube, one is to the
diode array and othe rto a sensor?
Under the preset board for sensing laser on? and diode power?
With no disc in place to reflect power back in?
To rebuild, support th epcb to the Al shell or the 16w
kaptan/ foil ribbon will split, which this did so abandoned
The 5 traces to the 2x presets are for the 8 pin opto device behind the presets as well,
so associated together for power monitoring and control?
Explaining the U61 situation of dropping out of operation
if no disc is in place
Panasonic NV D80 vcr smps problem
Failure to start up.
After removing the power unit and checking this and that
but not actually replacing anything then refitting in the chassis
the m/c would work a couple of times then fail to start,never
starting again until repeating the same abortive procedure.
There was some sort of "memory"
effect going on.
The 47uf,63V electrolytic was leaky on the supply side daughter board,
On removal this cap showed 5micro-amp leakage on 9V and
on an ac capacitance meter showed about 2microfarad.
Panasonic NV D80 vcr.
Non ingress of cassette.
Large toothed pulley driven by toothed belt had
fallen off spindle as retaining circlip had failed.
For the archives the power supply voltages should be
1 Gnd,2 Gnd,3 Reg 5.1V,4 Reg 6V,5 12.3V,6 power off 'L' ,7 Unreg 14V,
8 Reg 45V, 9 Motor Gnd, 10 Reg -30V, 11 Heater +,12 Heater - ,13 non-sw 12V,
14 LED Gnd, 15 Unreg 8V.
Also in p.s. R08 and R35 should be 0.39 ohm.
Panasonic NV D80 vcr
Tuner memory failure
After an hour or so of m/c being powered up,on
changing channel the tuning voltage would be at top
of range (32v) and m/c useless until cold again.
Squirting freezer spray on the EEprom MN1220
under the display bezel restored function again (temporarily).
As could not find a DIP version of the MN1220
put a zener in the line to pin 9 of the IC to
drop the voltage from a bit to cure the temp
build up and restore normal function read and
write .Also formed a
ventillation hole in body under this area.
Panasonic NV D80.
Behind the front panel the eject button had partially
sheared and jammed the click switch in active
position.Also the memory capacitor had failed,to
avoid total dissassembly of the innards,desoldered
the errant cap in situ and allowed to drop out and
replaced with a new cap mounted off the board wired to the
Panasonic NVJ35B VCR
No tape igress or other mechanical functions
The owner had forcibly removed jammed tape and had made the mechanical gear
trains jump teeth and go out of sequence.Turning the large pulley manually
along with manually triggering the solenoid mechanism with the inner right
part of the cassette housing unscrewed and raised to disengage the metal
rack drive allowed resequencing.
Coincidence of both radial tape guide arms in extended position at end of
drive sequence coinciding with cassette in down and locked position
Panasonic NV L25 vcr
Broken cassette carrier mechanism (heavy handed kid)
The 2 flimsey restraint/guide pieces of plastic that hold the
two sliding metal actuators had sheared.For alignment
concerning engagement of quadrant gear on the cassette
hold-in mechanism the last notch of the metal rack engages
with the second to last tooth of this quadrant cog.As far as i
can see the hook like mechanism that latches the cassette-
in mechanism down to the deck is only for assembly .
Find the neutral position of the main drive pulley system by
rotating by finger the large drive pulley and operating
the solenoid or the connected pin on the underside.
Latch the carrier in the down position to reassemble inside the
chassis (with no cassette-defeat the protections) then with
a small screwdriver from the top de-latch this hook.
The steel drive cog that engages with the cassette carrier rack when at
the cassette-in position the index notch should be furthest from
centre of machine.
One of the 2 broken guides needs to be replaced with metal ,
hooked through the body of the carrier housing.Metal about 1mmx15mmx5mm
shaped as half a fylfot/gammadion cross (that foxed you)-
shaped like a riser and tread and riser in section of a staircase
and glued in.The other larger piece can be satisfactorily bodged
with a piece of scrap plastic hot melt glued to shear point
and part of the carrier housing but avoid interfering with the end of tape
Unfortunately the whole gear assembley (about 27 cogs) had jumped teeth due to
the plate that holds the capstan-flywheel brake and toothed belt tensioner
had probably stretched between the anchor points of the spindles for the
main mode-change locus-cam and the spindle of the sun and planet assembly.
With a couple of pliers bend top and bottom at the joggle in this plate to
reduce the separation of the anchor points.
For realignment of the 4 separate gear trains from the main drive.
1 Sliding tape guide pins.With tape retracted into cassette the 2 index holes
in the L and R drive cogs shold be adjascent and the hole in the
steel quadrant drive also at point nearest the cog.
2 Main drive leading to the mode switch.
Remove the cassette carrier but retain ribbon connection and
set the carrier in the cassette-in position.
On power up the m/c should settle into the neutral position
with the radial guide arm extended.With the power off with finger
in the large belt drive pulley turn back so the cassette carrier
metal cog rotates back to the cassette-in position.
Continue rotating back the main drive and trigger the solenoid twice
should bring the gear train to the neutral position.
On the mode switch there are two small pegs the smaller should be opposite the
dimple next to the double V on the static part of the switch.This small peg lies
over the switch contact.When all is up and running properly the small pin hole in
the mode switch cog should lie opposite the similar hole in the helter-skelter
cam in the m/c neutral position at 6 o'clock position.
If the m/c ,front panel towards you, then there are 5 connections
to the mode switch 1...5 with pin 5 at 12 o'clock position of mode switch.
Then index position is just after 3 o'clock.
Pin 2 is the common .The index position is the same after power-up
initializing and REW and FF and STOP tape functions.
Play position of mode switch is about 10 o'clock.
Concerning the pinch wheel assembly.
Set m/c in position just prior to cassette eject.The small hole in the mode switch cog
opposite the index mark is the point
of most retraction of radial tape guide.Reassemble helter-skelter
cam to match this position.Add the bobbin sub-assembly and the top double anchor.
For adjusting relative position of mode switch cog it is only
necessary to remove the top anchor and bobbin assembly and move
the helter skelter upwards after depressing the paul locking on the spindle.
If the m/c seems temperamental re gear alignment but is nearly fuctioning
sometimes then with torsional finger pressure on the mode switch cog
try forcing clockwise or anticlockwise and if remaining faults
disappear then move the mode switch cog around one tooth.
3 The pin connecting the larger radial tape guide arm and the cog system had
sheared.This arm onlly moves between initial tape-in and the m/c netral
postion and m/c worked satisfactoraly permanently set in retracted position.
It probably avoids a partial slack tape when retracting back into cassette.
For the archives p.s. voltages 1to12 with m/c powered up in neutral are:-
The mechanical deck described above is the same as used in Grundig VS 520
Panasonic NV SD200 VCR
Jammed tape and noise like slipping belt on mode switch run-away drive despite no
slippable bands in this type of deck.
First problem the mode switch was stuck in an undefined mode-perhaps
"play" +"rewind" but pinch-wheel disengaged.Powering on and off
would not help and to remove the mechanical deck from the m/c it
is necessary to remove one fixing screw under the cassette.To remove
cassette disengage hold-down arms and springs that engage with pegs
protruding from each side of cassette carrier.Then unloop tape from
guides to give enough slack tape to remove the cassette (tape still
trapped behind pinch-wheel).To remove mode switch remove one screw and
a plastic push-fit stand-off released from cassette side.
The original problem is the nasty piece of 8x4mm plastic on the mode switch
motor spindle.This in turn engages with the worm-wheel with axial latitude
to avoid jamming that is possible with high back-torque worm-drives and had split.
There is not enough material to try gluing and reinforcing so engineered a
replacement fashioned from a piece of brass rod of right diameter and ground down
with a Dremmel and ball-mill to give 2 prongs .Drill to motor spindle diameter and
solder to this spindle by pre-heating with hot-air gun and then soldering.
Power supply voltages:-
Pye DV105 VCR
One head working other leading to vertical blind noise and distorted sound.
Unclipped the "head cleaner" arm and junked.
Manually cleaned with a piece of paper,power off. Replace torx cabinet screws
with ordinary self-taps
SAISHO VH1200 VCR
No tape transport
Radial tape-guide arm on take up side was jammed
against pinch wheel due to stel pin that
connects to cam system having dropped into
base of m/c due to retaining plastic having cracked
Saisho VR 1000
Intermittant failure to retract tape into cassette on stop.
Replaced the spring that holds the jockey-wheel against the spool drive
motor spindle with a stronger spring. It also cured a long term intermittant
squeal while in play/record modes.
Saisho VR1000 vcr
Moire fringing on picture.
The earthing spring connector was failing to make contact
with the bottom enclosure plate.
Saisho VR1000 vcr
Loud squeel from m/c.
The owner cured temporarily by propping up the m/c then moving to
a different position when it started again. Due to out of parallel between
axis of pinchwheel and capstan spindle.
Saisho VR 1000 vcr
Play function drops out after 6 seconds.
FF and REW ok,play ok for the 6 seconds then tape stops and retracts.
Dirt in the rotation sensor of the take-up spool so falsely registering
as lack of rotation.
Saisho VR1200HQ vcr
No system functions
Replace the power trannie next to the STK5332
with an uprated device (if TO220 then insulate
from the STK regulator.
Saisho VR3300X vcr
Intermittant pause of tape transport
I thought binding of the capstan drive spindle had caused
loading of the 3 phase driver chip M51782..Cooling the chip with freezer
spray temporarily cleared the fault.To remove this heavily
heatsinked IC melt solder around central IC tabs and with a small screwdriver
between IC and heatsink remove the aluminium then prize up the tabs then desolder
as a normal IC.Unfortunately the same fault emerged;it was in fact worn
bearing on the capstan spindle causing the trouble so replaced this and all fine.
Curiously on REW and FF the spindle must have centred itself with no loading
of the IC,only running hot in play (low speed).
Samsung NI 14N3, 2004,Combination TV,VCR, single tuner
Requiring frequent head cleaning because of noisy
picture despite less than 1 year old.
Noisy replay on this machine good played on another machine.
Lot of fine talc-like powder in the VCR section but the
TV section almost spotless. Blew out unit.
Is this another supposed head cleaner gizmo?
Next to the supply side tape guide, on the curved-slotted
pathway, is a brass cylinder that touches the active
side of the tape , rotating with its motion.
I removed it, assuming it is some idea of a duff tape cleaner.
I tend to always remove those nasty foam supposed
head cleaners, on placcy arms that engage with the spinning heads ,
as seem to CAUSE dirty heads. To remove just unscrew one edge
of the screen between TV and VCR it is not positively held
on the pin.
I've never seen one of these gizmos in any VCR.
It is engaged with the tape in all modes,
Play, Play+FF,Play + REW, FF,REW,Express FF & Express REW.
There was a black smudge on this cylinder thingy before I
removed it. Don't know what it was ,didn't blow off or
lift off with a magnet. Probably something vaguely greasy
that came off by wiping with finger.
Since removing this brass cylinder and playing/recording on
different tapes the noisy picture has not returned - so far anyway.
On first receiving and running first time of rec then play
was a good picture, second R+P then noisy picture.
Removed this gizmo and repeated 12 times on 6 different tapes
and no return so far.
Black smudge was something like the way photocopier toner sticks
to surfaces because of the slight surfactant/silicone content. I've never
seen such a smudge on impedance rollers before so can only
assume happened because brass rather than plastic.
Ribbon connector dc voltages going to VCR section in play mode
Samsung SI7220 vcr
Excessive braking on the take-up reel
Samsung SI 7220 VCR
Reported intermittent loss of picture (to black level) and sound
to mute not noise,on trying also on E-E.
The tuner seemed rather warm,so removed cover and on
digital probing (prodding with my fingers) the TD 6358N
would appear to have had a dry joint.Removed tuner,re-did
all nearby solder joints and left the cover off the tuner
(no ill effects) in case of thermal induced fault.
Samsung SI 7220 vcr
No transport functions but motor noise.
Usual old slipping belt on the mode change
mechanism drive motor pulley.The point of
mentioning is that it is possible to remove old
belt,clean pulleys and replace new belt through
the bottom access without taking the deck apart.
There is just clearance around the mode mechanism to
stretch the belt and pull through the gap.
Use a small S hook (like a shepherds crook),using the
outermost bend to hold the belt while looping onto
the recessed motor pulley.
Samsung SV 627B vcr
Intermittently running slow but never dropping out.
Never occured with me checking it. Inside was a label
floating around under the deck unstuck from a cassette and I replaced the
pinch wheel as glossy. Machine didn't return from owner.
Renoving ribbon cables to heads ,wrap cloth tape strips
around ends before re-remaking connections to beef up.
If removing deck make sure 2 idents are lined up on mode switch
on pcb and 2 idents on cross slide rack under the deck before reassembly. Remove front
panel to get glimpse under deck when reassembling to align
these 2 sections together then mate the connectors at the rear of the deck.
If no spool drive on reassembly then probably these 2 mode sw sections
Some DC voltages on m/c in neutral.
Measured on large inductors L401 5V,801 12V,802 5V,603 12V
Samsung VI 730 VCR
No sound on record but plays prerecorded tapes
No trace of bias oscillation at record head or the full-erase head.The bias
oscillator is housed totally in the
can that is the size of a dual IF can,the 2SC1318
in this assembly was duff.
Samsung VX972 video cassette player ,12V
Slipping belt on mode switch drive.
Another of those situations where just a couple of
mm cut in the chassis the belt could be changed without
taking to pieces. Remove the 5 large screws to release the whole
mechanical deck. Mark all cogs and levers of the mode switch
mechanism before taking apart. On reassembly make
sure the back tension brake radial arm is on the correct
side of the trailing tape guide traveller.
Sanyo VHR3100 VCR
Poor and absent rewind
Replace the tyre on the final drive pulley (surface breaking down) on the jockey
assembly or remove,turn inside out and replace.Notes: the earthing straps
to the ground in area of head drum and cassette carriage are necessary or the
m/c will stay in standby mode.Also beware of broken tracks/contacts
on the head drom motor driver board because of proximity to flimsy deflectable baseplate
to the m/c giving intermittent drum rotation and play cessation and unlacing.
Schneider DVD 857, 2003
no response to front controls
ribbon stuck to bottom of chassis .
Undo 4 nuts and bolts , upturn and manually move tray out .
functions via R/C ok, then pressed "key lock" and front panel
control returns, no red key icon or anything on the CCD dispay, in this mode.
To replace the drawer drive band, with the drawer in the eject posistion , the band
is easily accessible.
To de-elevate plater/eject drawer, with the deck removed from chassis and turned upside down
there is a black slide/cam plate across the width that can be pushed/pulled by hand.
Sharp VC 651 is much like vcr Philips VR 6542
except for front panel+ tuner and the rear panel.
Sharp VC 8381H VCR
Failure to fully eject the cassette
Positional latitude for cassette-in is provided by
the springs mounted inside the dual concentric large cogs either side of the cassette
carriage,similar latitude for cassette-ejected is provided by the smaller dual
concentric cog assembly on the left near the motor.Replace one of the missing a
nchors on the cog with a metal pin bonded in and find and/or replace the spring
Sharp VC A502 vcr
Idler tyre breaking up and pulley on shaft of mode switch
motor spindle loose.Dismount motor by removing whole
deck section and deforming the spindle with mole-grips
(lock-jaw pliers) or double-action pincers to give purchase on the plastic pulley.
The ribbon cable to the cassette carrier is disconnected by
pushing on the socket housing.
Sharp VC-D805H VCR
Erratic drum and capstan lock followed by tape
Poor contact in one of the connectors on the ribbon
connecting the capstan driver board to the drum driver
Sharp VC MH711 vcr
For tracking adjustment press up or down channel
or both together to engage auto tracking.
Super fast REW or FF on VCR, problem, Sharp VC MH711
Will rewind 90 percent of a 4 hour tape in 40 seconds or so and 20s for the
What would allow the motor to run at full tilt ? clunking loudly to a stop.
If it was not for the "crumple zone" of tape pulled into the gap between
tape roll and spool disc then the tape would have snapped at the end of
What to look for in the way of a problem ?
System presumably monitors ratio of supply and take up rotation rate and
direction of wind and unlaces tape from drum and over 3 stepped increases ,
sets at maximum revs until the ratio goes over a preset value and drops by
stages to a slow rate until end of tape sensor acts.
It should recognize getting close to the end of tape and start reducing the
power / revs of the motor but it once continued at full revs to the end of
tape. Been monitored by human since it happened. It does not speed up if
human stops the motor close to end, goes normal speed to the normal clear
tape auto stop if REW is engaged again ofter pre-emptive stop.
I'm thinking a break in the line from the fast , take up spool sensor, so
less pulses coming from it and false interpretation.
Taking the cover off and visually monitoring the amount of tape left, the
VCR works fine FF or REW . Going up to express rate and back down , in
either direction, as long as there is plenty of tape run..
It just once , so far, spectacularly failed.
Left as is with human monitoring. Owner always rewinds and
then FF 2 seconds to avoid normal tape problems.
If the auto-rev-reduction fails then the ruffled
tape going around the cassette guides will slow it anyway.
Easy to undo the 5 screws, take tohe cassette top off
and unwrap the crumpled tape.
Sharp VC MH711, 1998
Ghost in the machine sometimes osd menu pops up , sometimes all
segments of CCD show .
Took apart to see if reseating main deck helped as preceeded by vcr being knocked.
Beware high V on 400V cap long after unplugging.
While in there changed the clutch spring as sluggish tape REW.
Remove plastic circlip to get to clutch assembly.
Pull white disc forwards to release rear plastic gripper. Remount with
white gears in right place to slide , over black splines, compressing springs.
Original spring was 60gm to compress in state as used. Replacement about 600 gram was too much.
Caused loss of treble, loss of all sound and rucking of tape at pulled side of pinch wheel.
Found a 300 gm one.
Inspected mode sw motor for goodness of fit of worm drive on spindle. Unclip 2 main white
clips and then one small joggling around at the active end. When released nothing holds the pulleys in, 6V
is enough to test the motor /cog fit
Cam alignment seems to be a ruler laying against spindle point and then 104mm
away a similar brass point then when the mode is as ejected tape then small hole in the black body of cam lied under the ruler edge.
On the cassette lifting PTO number 5 of 6 radial spokes of the white half cog lies under a half
hole on cam periphery. May have reassembled with opne tooth out as although all functions ok,
inputting a tape was only possible by pressing standby button at same time.
Sharp VC MH711 vcr, 1998
Got a sharp knock , not dropped
Inability to set in timer mode, despite tape in place and programming set, then select
timer mode and the clock flashes in the display.
Beware of possibility of HV on th e400V cap, R discharge between anode of D912 and
cathode of D910. Vertical leaf sw for tape in sensing was dirty, unclip clear plastic
cover. Resoldered IC701 with solder paste, ridding surplus with toothpick and then meths wash.
No pin to pin connections and none after soldering. Same no pin to pin connections with IC801
Removed smoked platic cover to display - nothing wrong with brilliance , just
designer;s foible - much more visible display afterwards
Sharp FSP 2 , IC701, 1X1420gen4
IC 801 MN12510F
IC 702, BA6977S
IC651 BA7755A for audio head
IC1701 MSP 3415D-PO-A1
reset IC VHIPST600K replaced with T520C as a lot of green corrossion under its body
and on its legs as though some electrolyte inside this TO92
Sharp VC MH713 vcr, looks similar to VC MH711, differnet audio/control head assembly.
Squeel in use, cleaned the supply side band brake.
If FF or REW not working check whether VCR or not mode is in the display, via MENU
Remove the foam "dirtier" arm from touching the drum.
No picture or sound on pre-recorded tape play.
Someone had played with the sound/control head. Putting in playx2 or FF+view would give
a picture for a while. Adjust those mounting sc until a picture in play mode and then
adjust for sound , any and then high f.
Check for jitter in playx2 and FF+view as well as good sound.
Also record as LP and check play back and funtioniong of the counter.
If no picture then press the MENU button
Both video plus version of R/C GA134WJ and GO247AJ work on this vcr, "universal" R/C
main deck feunctions , MENU but no set to TIMER and cannot remove DISPLAY data from
Sharp VCT310 vcr
Disrupted gear train,the cam follower pin had jumped for the mechanism that drives
the radial arm that pivots under the audio head. Bad design on the back tension arm
side. With wear of the felt on the brake this arm can rotate beyond the point where
the tape guide on the left hand slide arc intercepts this arm when in tape retract mode.
The slide being worn could also be to blame but either way the two can jam together
and then the cam follower jumped. Unscrew the bolt holding the upper and lower parts of
the slide guide post and fix a modified solder tag to extend the interception surface
and another modified solder tag screwed to a convenient hole in the base pannel to
stop the back tension arm moving too far leftwards. Properly set gear trains and levers
are in stop mode:
Drum lead in / out slide guides notch and peg coincident and small hole and small
Small hole in triangular portion of locus cam coincident with a small hole in
In cassette ejected mode the radial arm cam follower should be in the outer slot and
when in cassette-in (stop) mode should be in inner slot.
PS DC voltages on 10 way connector 12,0,9.5,4.7,49,AC,0,-25,-25,0
Sony SLV 715 vcr
Tape jamming on take-up side.
The pivot of the take-up side radial arm tape-guide was sticking so
would not retract with the actuator cam.The lubricant between
gudgeon and pin had deteriorated.Disassemble and clean pin and
"ream out" the gudgeon with a wooden matchstick and reassemble
Sony SLV 715 vcr
No functions ,no response to on/standby switch.
No 12V rail ,in the ps the supply to the IC203 regulator
had a poor joint.
Sony SLV 715UB vcr.
No front panel display
There was no -30V (printed on pcb as 30V)or luminescent display heater
voltage.In the ps C203 1500uF,25V had leaked electrolyte
which had corroded /charred a hole through the pcb in an
area with no components,between traces for the positive side
of this cap and one of the display heater lines in the area
(but not involving)C214.The electrolyte had continued its
destruction on the track side of the board,20mm of one trace had electrochemically disappeared.
Made good all tracks,clean all surfaces of electrolyte
an abraided away the charring with a dremel (charred pcb resin material is conductive)
,replaced the C203 cap and an open circuit R202 150 ohm,.5W of the
For the archives CN412 power rails should be
1 GND,2 Unsw 5.8V,3 -30V,4 heater 0,5 heater 3.2V,6 nc,7 power on,
8 Gnd,9 12V motor,10 12V motor
Telefunken 1950 VCR
Tape jam,no transport or eject
Worn brush fingers in the lacing motor
leading to short and knocking out 1.5 Amp
pcb fuse marked F15A that feeds the M54544
Toshiba D-R160 DVD recorder, 2009
Probably a generic DVD mechanics question.
One motor for sledge , optical deck up/down and drawer open/close.
Transported by courier and would not accept DVD when received at the other
end - box must have been jarred. Try pushing the drawer in and it knocks the
spinner motor assembly which is in the up / play position. The repair manual
only says replace the DVD unit in this situation. Removed this sub-unit,
N78E3CBN no info found. Delatched the drawer and removed , turned cogs and Z
track etc until only the drawer motion cog moved with no movement of sledge
or Z-rack and reintroduced the tray and optical unit in eject position. Is
this a standard approach to this symptom for direct coupled ie no drive
bands? Seems to work now - but just temporary or fixed proper for normal
use? No cracked /broken cogs/racks or free-floating pieces found.
The drawer seems very flimsey, I assume it flexed and a guide peg got on the
wrong side of a rail.
looks as though the optical deck jolted from eject to play position , in
transit. Then first use after , nasty scraping noise of drawer sliding over
This motor looked more like a button rather than motor and platter, it was so
At one time with CD decks there were transit screws like vinyl decks. Did
some change do away with that requirement for CD and presumably DVD decks.
Was it just a change of suspension? originally great floppy springy things
and then somone realised you only needed to stop audio feedback, so in came
little soft grommet like things .
Pull out 2 wide ribbons , reinforce to reconnect
remove 4 sc of mechanical deck
Then lift 2 clips at front and bring deck over pc pulling drawer through the front panel slot/
Leaves DVD mechanical deck and first board connected
Release 2 side pawls to drawer and it will slide out
REW but not reliable Play or FF
Slip clutch problem. Remove cassette carrier. Prize off the top
cap on the spindle of the slip clutch after disconnecting the hold on spring
carefully not to let the "clutch -plate " spring go flying. Replace the pulley
tyre ,stretch this clutch plate spring a bit and before fitting the other spring
reduce its length a bit.
ps DC voltages from transformer end
19,.7 // 9,0,12 // 9,NC,11.7,0,13.5,0,12,2.8 // 11,-21,-21,0,0,3.9,1.8,-60
Toshiba V631 VCR, 1999
Would play pre-recorded tapes only - very snowy picture.
Comet tails on picture noise and 2 of the 4 heads were
chipped in central gap section.
Would only play as AUTO with record switch open from
cassette ingress otherwise no deck
functions - no play,REW,FF or stop or eject. Stop engaged
by pressing REC and eject by pressing Standby.
Central source ,End and Start sensors OK and
3 contacts of mode switch seemed ok. Didn't proceed
Toshiba V703B VCR
Modulator has no UHF o/p including test "card"
Heavy handed adjustment of the UHF channel change had
broken the retaining clip for the moving vane spindle of the
tuning capacitor allowing MB to short to earth knocking
out the "TO-92" packaged pcb fuse Z681 marked N5 ,replaced
with a 500mA fuse.
True tape counters
Beware of apparent non-function of tape counter
where the tape is brand new with no control pulses
recorded on the tape.
VCR and TV tuners with broken co-ax connector plugs
The mechanical strength of the junctions between
these connectors and the tin plate housing of the tuner
get flimsier and flimsier with later models.As the pin on the connector is
usually soldered direct to the tuner/modulator pcb it is
usually impossible to make a good job of reassembling.Remove
the covers to the tuner and strip back some RG178/BU or
similar 3mm coax (to be able to curve around to good fixing pad inside the tuner)and s
older to a good bond point on the pcb.
Find a good earth point for the shield and make sure the cable is well anchored
to the external casing with cable ties etc
and solder on the relevant coax plug or socket to be back
in operation.Often possible to do all this without removing
the tuner /modulator etc from the machine.
Televisions and Monitors
Acer XD1150, DSV0602, 2006
Lamp lights for 2 seconds , no fans , then flashing red LED.
Left and under plate comes away as one.
Careful removing top cover as thin ribbon to top sw panel, lift
brown flap of ZIF closure at the pcb.
4 side sc and 2x pcb sc to remove sides and rear cover metal.
2 wires at main ps , s/c, so lamp safet sw ok.
Beware colour wheel is very exposed
3sc on fan cover + 2 pcb pull off at main header on top main pcb
to remove fan with its cover.
Fan ok running on 6V so replaced.
12V fan supply from U44 on underside of main pcb, it
supplies p2+3 &5 of ps 16w
Blue for 2 sec until lamp is lit , then
both LAMP and TEMP once, then revert to slow RED Power LED flashing.
Remove the IEC for a few minutes and on reconnection the
axial fan spins for a bit.
To check central blower fan, dot of white tipex on the centre boss
and view over the colour-wheel.
Spins at first connection to power like the othe rfan, but does not
spin on pressing power button.
U11 supplies blower fan
Colour wheel ribbon conductors were showing a bit as though not pushed home, but were. Someone had been in there before
and replaced this ribbon upside down, blue strip should
So that LED error sequence is for incorrect colour wheel function/loss.
Reason for someone going in htere in the first place not found,
and ran for 2x hour long test runs
Akai CT2870E TV
Select preset number
+ or - for channel
Press M briefly twice to store in memory
Akai CT 2870E TV (dead)
Dry joint to the point of being spark erroded to no contact on
one of the pins of the L part of the LC mains filter next to
the mains switch. Wired in and resoldered all similar points.
Soon after repairing the 2A fuse failed probably due to
excessive loading by repeated surge current when the
dry joint was failing.
Akai IRS FX 5652 1323 TV remote control
Intermittant failure to operate until one ofa few
keys in one area were pressed.
Remade the solder thru board solder bridges
in area that 'cured' the problem.
August LP170B presenter
for advancing powerpoint presentations , receiver in USB socket
and a laser pointer with the transmitter pen.
No laser output, 2V,3V, 1.6V on laser pins
Q1 trace down to laser end , 3V
Aside as thought a new laser was required, a Hualei half pen size
cheapo, spring brass disc, lens smooth face to laser and focus by turning
outer closure. 90R SM white ceramic dropper supply of 4.5V
1.9V over 90R for 4.3V , 20mA DC
2.2V over 90R for 4.5V ,24m\A supply
so about 28R for 3V supply
Broke the fine wire contact to the laser die. From smallest slide sw,
grabbed a phosphor bronze contact and cut it down . Soldered to
top of laser mount to make spring contact to the laser.
Removed the laser and tried on a bench ps, droppered from 2.1V
and laser worked. Returned to the pointer wth cut trace to the
feedback photodiode and worked. As long as not left lit for too long
probably be ok. 4.7K on the pcb to the photodiode, added
another 4.7K and still held off. Would work with 100K added there.
Thick rubber band , pierced, slid over the barrel to
avoid accidently going into B mode, whatever that is for,
prints letter b on notepad or a powerpoint slide, or exits a powepoint .
Compaq FP5315 , 2003, also called Pavillion
conventional solder not PbF
see also tips on hot electrolytics
2x2 sc , plectrum down the sides , remove stand
undo the 2 outlier connector headers
undo sc to IEC clamp
sc to tinplate under tape under the central header
overheated ZD904 and gone ohmic measured cold as 26R
TL1451AC, 35AC9HT , Si4411 probably and K2996 driver
Hot spot C213 220nF,160V solder ring crack but cap ok,
D910 and D912 both ends poor solder also C5706 and wire jumper
blown 2A SMD fuse connected to ZD904.
Added vent holes on underside of screening and enlarged every other hole at the top
and drilled out some of the holes at the top of the plastic casing
Inverter 244R o/p to i/p
o/p in series via feedback Tr measuring 385R split 192/192.
So both tubes run in series from one inverter so this mid point must check
for both being triggered and correct power use
Primary .1,(210/195R) , .1R, 0, .1, (195/210R)
holes thru pcb
2 x 7.5V ,1W zeners, same batch, in parallel with insulated
twisted leads to replace ohmic zener
Q201 , Q202 topmarked 86 and 76 no DVM dide test good as digital Tr
76 DTA144 pnp and 86 DTC144 npn
removed Q209 and Q210 as shorted and ohmic
Used 2xTIP41C with cut down heatsinks
fires up with "check video cable)" box in screen. Switched off to add
a signal and restarting resumed to flashing green LED and click in speaker
Cs reasonable ESR
Would work from cold but bad caps includeing shrunk covers
Replaced C930 , C924 no change
replaced C929, C926 and C925 with 2x 470uF 16V in parallel ,
upside down and wired in was space.
Unnecessarily changed th e7.5V zeners which were overheating due to these caps ,
with 3x paired up 2V4 zeners and 1R to monitor current and silicone rubber shroud
In working order o/p 1 to 12
12.4V was 13.5V and 14.5V with duff caps in place
leave abit of a gap lower end of backing to insert bracket and then close down the plastic
outer rear case. Didn't think to measure V and I on my 1R when working well but must be less than
.14amp for 12.5V at o/p
measured mains power consumption at 15W mainly white screen
No control over brightness, 0 to 100 percent only gave about 90 to 100 percent.
Replaced 3 remaining Hermei caps C201,932,933 201 to 200uF 25V but no change with brightness.
p2 of 12 way conn varied .1 to 4.6V on OSD brightness setting.
p2 via 75K and 18K to gnd of dual PCM IC, 1415AC, p6 via 47K to the common point.
1K over the 18K brought brightness down to 30 percent or so, no OSD control on that.
Probably wrong replacement pair of transistor drivers.
Instead of replacing again, settled on 470R and 100K log pot set into the casing for
With about 30 percent brightness, resting thermometer over top of casing for room temp of
16 deg C, 24 C over original casing and 30 over the butchered section.
Not known what it was before all these holes but at least a difference.
Compaq 420S colour monitor
Intermittent slow fading of image, slowly varying
brightness or failure to bright-up at switch on
until the casing is knocked. Also occassional
sizzle discharge noise accompanied by slight width
reduction and loss of resolution in parts of output.
The slow fade was the give away. Poor solder
joint on CRT connector to back pcb in the heater lines.
PCB holes were too large for the pins, so cut small loose
spirals of tinned copper wire and soldered around pins to pcb pads
on all pins. Also the solder to the 3 colour driver power
trannies looked inadequate.
CTX 1565D Computer Colour Monitor
Black screen but on-lamp ok and yoke crackle on switch on.
Bad solder joints on 5 of the 8 pins between CRT socket
and CRT mounted PCB. The pcb holes for this CRT base socket were
obviously made for batch of sockets with much larger pins so
leaving too large a gap to be
filled with solder. Slid a small brass washer down each of the pins
and bedded into the solder pads. The owner said the monitor had not
been dropped or
the rear pannel pushed or knocked so there is probably a lot of
monitors of this model with this fault. Unsolder the earth straps and the 4 earth
bond point fixings to the pcb to remove the tinplate screening.
Generic TV r/c fault
Total failure of r/c
To mount the ceramic resonators they usually need to have their
legs bent but this stresses the plastic that seals the 2 parts
of the housing and eventually can force the 2 parts apart and the
resonator is no longer held by spring action.
Goodmans LCD TV schema try looking under Toshiba (different model number ) and vice versa
Goodmans LD1575D LCD TV, 2008, no PbF marking but probably pbf solder
LCD TV SMPS LV C looks like vented electrolyte, seemingly, evenly over the whole inside
of the TV and over all the boards. Leave in place? , wash off as best can do
with methylated spirits or something?. Otherwise the TV works of for half a
second before it decides not enough power available.
This looks like cod liver oil in colour and
texture , with no hint of corrossion on any surface it is laying on.
pH of 5 this yellow oily stuff, not conductive as far as paper soaked in it
, 0.5mm probe tip separation and >30M
Looks as though all this elecrolyte was a bum steer, leaking yes, but not
reason for failure. Will have to go in deeper . Puting in the pair of
Fairchild backlight inverter IC numbers FAN7314 and FDS8962C into Google - led to a French
repair forum with a pic of the ps of a Toshiba
and then schematic
which seems to be almost the same as this Goodmans LD1575D
I reckon the "electrolyte" is kitchen fat. I saked the owner if used in a
kitchen and he said no. But it was used in a caravan which was used for
cooking in. So tried tasting somer on the tip of my tongue and swilling off
soon after. No sweetness , if anything cooking fat taste. Checked my taste
buds with some glycerin and distinct sweet taste. No problems with cracked
tops , ESR or capacity with the removed and then replaced caps.
Cleaned the main board with meths , dried for an hour or so and same
Cleaned the ps and inverter board , dried for only 10 minutes and ps went
Dried with hot air and pic flashed up for 1/10 sec or so. Concentrated
heating on the 6 inch long strip of isolated track along the pcb edge to one
of the backlight connectors and now about 1/3 second. Previously about 1
second when I eceived the TV.
Agrees with the owner in that when this problem fisst appeared someone
technical decided to run a hair drier in the top and it ran perfectly on for
a week or so.
Yesterday I picked up at a hamfest a range of 1G to 1T ohm resistors
coincidently. I reckon its a glass fibre osmotic problem and HV , like I've
seen on scope tripler/quint boards over the years.
Will try cutting those long lines and HV cable bridged there instead.
There must be another control line back from the inverter protect system,
back to the main control and then back out to the backlight control to the
If this is the failure mechanism then presumably a generic problem for this
sort of long pcb trace to a backlight connector and then cable proper.
Will have to see if there are reports of this problem and delve into
backlight inverter control a bit more.
Again generally, comparing a 15 inch LCD TV to 30 inch. Obviously the
inverter and lights would be x4 power but would anything else in the
electronics be much different?
problem was with the 3 small 3Kv rated blue ceramic caps at the inverter
output. First variation between this Goodmans and the Toshiba schematic.
Overlay numbering the same , 3KV the same but 3x 5pF on the Toshiba and 2
marked 12 presumably 12pF and one 4.7 presumably 4.7pF. Now to find dome 5KV
Disconnected them and powered up for long enough to get the display on for
some 20 seconds and blue screen with annunciators etc before I decided to
switch off and find some replacements.
found a number of 22pF 3KV and seriesed 5 of them up , 2x2 and 3 ,working fine now
Using principally a 8.5KV transformer and variac and 300M ohm 1000:1 divider
and meter and 3.3M and second DVM (acV) for current
original capacitors passed at
1.5KV , 9 to 10uA
2KV , 12 to 14uA
2.5KV , 15 to 17uA
BH25 Y0504 SMD
ADJ-B 0720 SMD
STR W6253 smps control , 5K6 R82 , 150R
Meths washed off ps and control board and flat ribbon ends and sockets
In passing I see a 1W 4M7 R under the SMPS Tx with
the opto coupler, never seen that before , connecting mains side to
secondaries side. For dry countries and static build up on aerial.
Inverter Tx approcx 0R1 and 722R
TV will work with top panel sw board disconnected but on reconnecting yellow to white,
beware of latching on/off sw.
Control board fuses F600 1, F603 4, F604 1
Vdc 3.3,1.8,8,5,12 on links near PL600
J180 BKL on/off
33V on J105 near tuner
J180 4.45V no picture , 0V pic on
PVcc Q607 p1,2,5,6
to L407 and L406 to PL403 to LCD display
Goodmans 206NS TV
Broken aerial input socket from tripped over trailing lead.
Remove the cover from the DT2 IV17D tuner,not the soldered one as there
are active returns to this cover. The central socket connector was loose but
instead of removing the soldered-in side abraid the small rf coil near the socket
and soler the conductor of a small piece of co-ax. Anchor the cable to the socket
and lead through a fashioned hole in the casing made under the socket hole.
For the records a suitable remote controller for this TV is R28B03.
Fading up and down level of brightness. If full range then 3 or 4 seconds to cycle
up or down. You could see the heater glow or not in those cyclings.
Eventually more off than on. Resolder the ring of pcb solder points at the
neck mounted pcb connector. Returned a few days later , and in a darkened
room and twisting/ bending the pcb , wearing thick gloves a fairly feint white low voltage
flash to one corner of the board. p5 of connector KB had a bad solder joint.
Infocus X2 image projector , 2004
Prior to that , not seen by me , only reported, there was a slightly
diagonal shadow across the very right hand edge of the screen, but still
being used like that.
The whole inside is covered in dust, including a very dusty reflector on the
The actual lamp is marked SHP58 Hg.
Discovered what the original problem was.
A rectangular tunnel between the colour wheel and the first lens, going
across the airpath of a fan. Made of silvered microscope slide glass , 4
pieces forming a rectangular tunnel about 30mm long 4 x 3.7mm internal bore
but the glue at the
corners failed with time or heat and the retaining spring pushes one edge in
a bit. With no silvering it would have absorbed more heat and probaly
rapidly failed. The mask , part of the clamp, was seriously black,
required grinding off, oven cleaner didn't touch the burnt on nicotine
tar or whatever.
Removing the spring and the tunnel collapsed, i'm surprised the fan air
didn't totally disloge them. If rebuilt, bend the vertical part of
the spring away so only held by the top should avoid similar movement.
I ditched them totally and replaced with some 0.4mm polished stainless
steel bend around into a square of 6mm outside dimension. Not
closed on the closure line, as my original fracture trying to close round.
Added 1mm of PTFE sheet on the bedding plane furthest from fan
and 0.5mm on the other plane as packing. For normal table-top
viewing id a shadow on the RH edge then obscuration of the light tunnel
at teh edge nearest the fan.
Putting a 50 watt floodlight mercury discharge lamp across the feed, either
way round produces nothing.
Putting a 110V neon across the feed there is a constant orange glow but no
initial flare - too short to be seen ?.
Putting a 150 watt spotlight discharge lamp there, nothing.
Even tried a 9 inch standard fluourescent tube in all orientations.
I've removed the ps board. The high power, LV, side of the lamp drive looks
ok, despite some varnish browning around some wide format 1K SM resistors.
But on the assumed HV section, the inverter size and type of hf transformer
4 segments to the wiring , 1 thick wire and 3 fine wire all the same fine
gauge, secondary coil 115 ohm,80mH, 0.22 mm primary wire.
2 pairs of SM 1K R on LV circuit and 2 seriesed 2.2K droppers on HV cct
No 12V is getting to the HV circuit of JRC 2901 quad
comparator/transistors/inverter transformer, because the Q501 B764 pass
transistor is held off by an isolated line from the video board. Taking the
output of the opto isolator low by 330R the HV circuit works.
I now know that a standard 50W floodlight mercury discharge lamp will work
on this circuit as will a 9 inch fluourescent tube and more diagnostically
380V neon runs fine but nothing without the HV circuit, 110 V very bright
and a low output without the HV , just the 62V power sustain voltage.
Main ps heatsink about -150V to ground, lamp o/p
approx 320 to 380V. ps board has a black standoff clip at centre.
inverter transformer 2 sets of 9 turns and 2 turns for feedback
2x D1865 for drive
7812A, K2917,K2647, 28061, K3568
100K, therm 18R NTC, 2x 0.47R, 2x 0.16R, therm 6.2R
th sw 100C ? Asahi US-602H / 100CH
The power switch has been toggled on
and the software has initialized or the pro-
jector has been powered off via the
The power switch has been toggled on
and the software is initializing, or the pro-
jector is powering down and the fans are
running to cool the lamp.
If the projector has been working but
exceeded its lamp life , reset the lamp hours by simultaneously pressing both
Volume buttons on the keypad and holding them for 10 seconds
solid red, An unidentifiable error. Lamp hours is just the in-use time
from last reset, not a function of lamp "wear".
2B47 IR receiver and amp pinning 3.3V,0,o/p
Osram Halopar 20, 68832 FL could supply the 60 mm
reflector. Remove front UV-block glass, held on only with
RTV silicone probably, and bass brass, burr out
the indents then tear away the brass. Try the 50 Watt
lamp set further back or try a short 150W 240V linear
floodlight lamp set in fire cement . Run off the mains via a
relay driven from the fan supply, Remember 2 types
of fire cement, try different positions with non setting
and finally use setting type. Water-glass maybe useful as well.
Had to grind away the ends of a mercury discharge lamps , wire
in and cover the rearward extension with a cover bolted
to the outside of the projector made of lamp fitting and Bulgin connector
high temperature plastic.
Results for this lamp , probably much like
the original in function . 300W overall consumption and 80V, 1.8 amp
lamp consumption, 145W sounds about right.
11,600 lux measuring with probe in the output from the lens. 50mm from the ring,
with no video and white "blank" screen . With A4 sheet illuminated and a photographic
light level meter, set for 100 ASA gave f4, 1/60 second reading. With ambient
of 17 deg C.
Using diode temperature probes, the parabolic reflector reached 85 deg C and 88 degree
after switch off. The light tunnel reached 46 deg C. The housing
between thermal switch and over the colourwheel reached 26 deg C
and 47 deg C at switch off.
Normal power use when warmed up 80V, 1.8 amp (measured with 0.1R
in line and hookup wire volts-drop). At first striking then drops to about 20V then increases
to 80V. If still hot when trying to power up
then starts at 1300 volts, drops to 64V with no lamp ,
470V down to 64V etc
until the fans have cooled it down enough.
Menu brightness change makes no
difference to lamp brightness. Changing to low power
setting does drop the light output, but volts
stay at 80 V so presumably current drops, used for lower
fan noise, very large screen display not required.
Blue screen , no signal, low power setting, with thermometer bulb placed
at lower left front where most heat is blown out took 10 minutes
to stabilise at 25 deg C over ambient.
To re-assemble top and bottom then push front over the lens ring
and then rear
InFocus LP420 video projector
Dumped but with what looks like a brand new lamp inside.
Rigging the lamp up to a 150W traditional dischage-lamp floodlight ballast with 50R dropper lamp lit with
1.1 amp mains draw.
I assume the projector failed and owner bought a new lamp but because a bit
of a work up changing them , left it inside. Lamp reset sequence makes no
difference, factory reset is via Menu screen. No response to the control
panel buttons as far as the same LEDs stay the same.
4 central LED lit (so it would seem but probably those 4 not lit and the other 9 and background light), sound of colour wheel spinning
4 way lamp connector 2 outer of 3 loop back as lockout safety
505-0261-04 / B0039 - 037 label
1F4 OA on lamp
1801225 on outer
With stripped down innards 9 orange LED and one green LED
Disconnect thermal sw or the lamp loopback then no LEDs
lamp current transformer , for lamp on monitoring ?
large black lump to lamp
47mH, .06 ohm / 1mH, <>0 ohm
inverter Tx 138R / .1R or so and cold testing of associated Qs had DVM D checks reasonable
431A1 76T C8D7 8 pin U407
2 mustard caps either side of the GDT
472M 3KV and 22M 3KV
2 seriesed D , 2V diode drop each
fed via D308 & 680uF 200V from mains via D302 bridge rect, its neg to Q500 and Q502
ps fixing sc are smaller than top pcb etc
To replace the top pcb with serious multiway headers each end , refit the colour wheel
kapton ribbon first.
RT 302 , 68R thermal sensor
RT 300 , 5.6R
Lamp current return via 4x 20R +2x 33R giving 9.1R
In situ, Q300 IRFP460 showed no state change with DVM D swapped lead test.
Remove bolt ,then heatsink to remove the TOP3 and then the DVM test worked
LV ps rails , 6 pins all ground 2x10 way header
5,5 (1.7A) ,3.3,3.3 (2.2A),12 (.8A),-,+/-9V ? , +/-9V ? , .25A each
000000, 5,-, A,2V
0 to frame, not the -mains neg
A transpired to be the lamp-on control line
opto pcb mid I called block 2 of pins
For an extension lead for this header , both plug and socket ends need grinding back to fit
available surrounding space.
p1 40K to p4 of lowest opto
p2 to Q103 = J32 via 10K
p3 to 7805
p2 to lowest opto from one of th e9V on the ps header
Block 2 DC
0,12,12.4, 6.6,0,0,2.5,.03,0,0 (to RT303),.1,-,0,5
pin1 of CS3865C double smps IC not soldered but not connected electrically perhaps
Rt 8.25K , Ct 2.2nF
output pins 7 and 10 to Block2 p5 and p7
p5 to Q304, p7 to D401
Desoldered th eopto board and temp mounted the other side with a corner brace
of hotmelt rod.
CS IC p14 is low so that side disabled. Out1 controls the LV ps rails
p14 from PICp7 via 1K R217
J32C 11.8,0,11.8V dc
Added 1K to R217 and fed from 12V to take p14 high would bring up 130KHz pulses to
Holding p14 high at switch on would disable th eCS
Confusion on the IRFP predriver areas due to Shottkey ? diode D111 on th eopto board ,
removed the 949 unnecessarily, breaking it in the process and replaced with a 2SB1202
Mini Centronics 36 way connector is type 1284 according to the manual
Trigger cct lines
p2 signal via R216 to p8 of PIC16C711
p4 Tx primary (feedback?) from Block2 p3
via diode from Block2 p1 and p3 of U100 34072 4ROV =MC34072?
With p14 taken high and a load of 2 20W 240V filament lamps as lamp load then
PIC p3 to U204 p2
p4 via 470R to p1
PIC p5 to U102 p4
CS p5 to U102 p5
U102 control input
Touching p6 the base of U201 with DVM probe and lamp lights.
Actually 100K from 5V to hold b H ,
Instead of filament lamps tried a CFL tube , just 2 of the 4 wires and 220R dropper.
Stays lit for 4 seconds after removing control line H
CFL and 2x 220R in series , lamp flickers
CFL and 2x 220R in parallel, lamp is too bright.
Returned to proper discharge lamp and full power useage, 1.2 amp and then 1.4 amp of
I've found the control
line for lamp-on and taking the downstream side of its opto coupler H , the
discharge lamp triggers and works fine . The 4 arrow windows probably lit from
other LED spill over and light rather than the 4 directly under the arrow windows.
The so-called service manual out there is no help as board-level replacement
"repair" only. All solder except the BGAs (as unseen) seem ok
All DC levels for the controllers seem ok. As this lamp-on control rapidly
goes to one of the 2 BGA proprietries labelled PENG and GUIN (have to be
called something) is that the end of any repair?
Signatures of assemblers? in the moulding of the underside inside of the plastic casing.
I will reassemble , now with a usable lamp , and see if anything turns up on
the video screen -no . Mains draw up to 1.5 amps and the variac supply , at 100 percent
, the 1.5 amp thermal cutout popped, probably just as well, 270W + 50W is less than that.
Replaced a poor ESR 100uF on the main control pcb but no uC control of the lamp.
Tried pressure on the GUIN and then PENG BGAs then other large ICs and no change.
Left at that , a known good 270W lamp, sans proprietory carrier, and parts mule
for some other projector in the future.
An aside on the very strong red epoxy? holding the SMD on the ps pcb
From Uffe Bærentsen "Have seen thermo curing glue used when SMD and leaded components are
used in a mix on a single sided PCB.
First step in production was to place all glue dots for smd.
Then place all SMD components and cure glue in a reflow oven.
Then place all leaded components and run through the wave solder machine.
In this proces both SMD and leaded components are soldered at the same
time and the glue is to assure that the SMD components do not fall of
during soldering. BTW the glue was red."
That sounds likely , a mixture of SMD and thru-board componentry on this
pcb. More glue than the dot that would normally be used just for holding SMD
in initial placement.
There was another feature of it I'd noted but dismised ,thinking it was just
due to thin film "dilution" of the colour. The exuding glue blobs around the
SMDs was bright red but the underlying thin material was more beige or peach
in colour - perhaps a side-effect of reflow oven and air exposure while
431A1 = ? TI TL431A1 , 8SOP version ?
ST MAL AVS1AC automatic voltage sw for 110/220V mains with AVS12CB triac
431A1 p 2,3,6,7 connected to p2 13K3// R412to give 10.55K in cct with fan to p1
p8 to LM317T, p1 to adj
p2 to FAN , Vin
p3 to 2x10 p9
Remove optical block - superglued screws?
grind a slot in one top sc as temp had no effect as far as I wished to take it.
Grind into plastic base to expose the hank on one underneath.
3x st/st clips hold output lens.
2 sc to release colour wheel section and 3 to release motor
LG KC 14P2S, TV + VCR , 2001
Broken mains sw
Remove case , VCR section 2 sc halfside , 2 at rear
Unplug 3x 2W conn
remove sound board
4 wire deflection coil conn
CRT EHT cap
vcr remove 3 conn, sound col, 2 ground bonds
remove top and bottom covers, short sc on side
4 deck sc
head ribbon and sound ribbon
alll to get to mains switch soldering
mains sw front face of false button about 1mm in front of pcb
unclip base pcb
redid solder and 2x mount points
4 pillars at rear of rear cover go over the TV chassis
Mitsubishi XD200U video projector, 2002
and some general discharge lamp observations
and later Towards the goal of a truly universal IR remote control
To remove cover over mains section there is a matching screw hidden
in central area under vertical wiring, more obvious if you remove the
outer lamp housing (original lamp and its housing missing)
102K across lamp supply and 0.42V "diode " test.
Small igniter Tr 7R, .3R//15R
The inkjet printer syndrome. Why pay for a new set of inks when a whole new
printer costs only a little more than that. Another thrown out video
projector, all that clever technology , just because a replacement bulb is
near enough the cost of a whole new projector. Anyone know of a sensible cost supply of
generic basic discharge small size lamps, intended for fanned operation?
This one, from the manual, was 210 watt NSH lamp, NSH = ??? = something
metalic nickel? (N ?) Sodium Halide ?
I have a few 150W "Power Star" discharge lamps that will physically just
about fit in the available space. Used in floodlights they don't need fans,
projector versions , as fan cooled , are physically smaller.
In a luminaire these 150W lamps take 0.55 to 0.6 amp at 240 V mains , in
use. Tried one on the projector with a 12V 50W filament halogen lamp in
series as a bit of a dropper.
939P901A10 by Ushio , 939P901O10 LAMP POWER UNIT in service manual,
uses Mitsubishi 34513M4150FP ,32 pin controller under the torroid.
With ammeter in mains supply , overall current (including SMPS and fans etc)
was 0.8 amp for 10 seconds or so and then jumped to 1.2 amp to full
brightness. I chickened out then and switched off. Where to go next ? Next
time I will dig out an old analogue ammeter (concerned about the presumed 4
or 5 KV starter/igniter) and place in series with the lamps.
No mention in the manual of this initial power step but it is the same with
my other projector so presumably generically normal. 6 wires connect to the
ballast/choke so presumably there is active switching of parts of the choke
which maybe selectively disableable and not monitored by the controller. I
can live with reduced light output, my other one has a preset option and is
set for half power maximum on normal user brightness control.
Existing floodlight using Osram HQ1 TS 150W W/NDL Power Star
Igniter ZRM 1,8ES/2 Ib max 2A 4 >4.5KVs
Ballast for 240V measured 3.9R , .35H, mains current .55 to .6 amp
Trying an 80W MBF/U in the floodlight with a 12V 50W halogen filament lamp in
series as a bit of a dropper, took .55A and 2V over lamp , very dim orange colour
Beware 2 5 pin connectors with XD200 ch5 at double fan board and another near the
main output chip.
supply sub unit is marked 370V dc to 210W self contained taking rectified mains ,
separate from electronics SMpS, and 3 opto-isolated control lines only. 2 in and 1 out
it seems, hopefully 2 power levels and lamp lit ok op.
2 25R in series then 240V current down from 1.5A down to 1.2A with tubular type lamp
after 30s (210W) 120V o/p
DC over each R dropped from 58V down to 44V so 2.3A , 77W each R so will try one 25R for 150W over lamp
Ballast and ignitor board 2x NEC SMD topmark 842 for upc842 and 14 pin C324G for upc324G so dual and quad
opamps . Adjusting the preset or changing the feedback chain may adjust the output, not tried yet.
J16 on optics section is empty.
Managed to leave the J14 ribbon in place on removing the side fans section and ripping the
socket from the polyamide ribbon. remove 4 screws and heatsink on projector chip.
ch(5) no connection
(3) .135V with main J3 cable connected
(2) To J3 ribbon , about pin 10 from outer edge and innermost line
Clear away side solder fixings, solder 5 pins , check continuity before hotmelt gluing to ribbon.
main ribbon is stiff due to mumetal shielding
LED red , fans , no lamp ignition , power alternate red /green if J14 not connected
Found a new lamp, cheap as surplus spare to a broken projector, for an Epson ELP 3500, MSCR150E3H / 1028301 . With exhaust fan moved sideways
to the outside and some of the Al webbing of the chassis managed to fit this lamp in place
with some silicone rubber pads as buffers to the chassis and rear mount fitted through the rear
high temp plastic casing with PTFE washers . Replaced non-mating lamp connector sections with
ceramic choc-bloc. Fitted 2 x 51R , 20W vitreous R in parallel to some phenolic tag board in
the space where the fan was.
This lamp, new 150W , 65V rated one, ignited with 70 percent mains , 50V over the vitreous droppers
and 1.2 amp disipation with 100 percent mains for a short test. Drape floppy silicone rubber
sheet over the optics section as adhoc cowling to get the fan air passing over the lens etc
Video projector without an IR R/C , originals or substitute not available
and cannot access the presets.
Flashing red and green status leds , the lamp cover closure sw is not a small
micro sw but leaf type, so no click noise.
To get to lamp sw , undo its sc, slacken 4 large sc at the base of the whole
optical unit . Release 2 sc either side of 1/2 circle lens shroud and turn it and
push inwards . Release metal clamp over large ribbon to the optics and
lift off. With a bit of space now you can see under the lens and ribbon to release the 2
plastic wiring loom clips and pull the lamp sw through.
Bn & Bk wires go to the small vertical board part of <>10 way connector at head of the front
IR receiver .
After all that Id misread the manual and red/gn flashing is "lamp life" so
just pressing Power + Auto Posistion for 3 sec a few times , then after 1 minute
at power up both lamps go green.
Putting duralinium top cove rin place and mains ELCB dropped out.
The outer most of the the 3 20W R was intruding too far and touching the metal of the top cover.
Top panel sw , 6 lines to p15
p1-, p2 +, p4 vid, p5 comp, p6 power, p7 auto pos
1.09 amps of mains with lamp on after 1 minute, 2 mins from power on.
60mA no lamp but fans on
With bulb of thermometer immediately outside the main fan outlet section registered
26 deg C over ambient when stablised, time not noted.
Found another lamp, physically smaller after removing the surrounding casing, for an
Infocus LP420 , lamp rated 270W
With 0.1R in series
85% mains, .9A mains, .41V over .1R
100% , ,8A mains, still 4.1 amp through bulb and 28V across bulb (measured after settling )
For setting focus position used an LDR to DVM set in the screen middle with
white projected. Used the original ceramic surround to the dichroic plus springs to the
Mitsu outer shell , and locked in place with RTV when maximum screen brightness found.
No looking directly at the lamp output of course.
After 5 mins of lamp on , measured 34.4V over the lamp.
If projector goes off totaslly, no lamp or LEDs or fan probably the bead
thermistor over the lamp housing causing overtemp protection.
Had forgotten about the hole I'd cut in the top of the PPE housing so mounting
plate directly headted. Placed some thick silicone sheeting over the hole clamped by the mounting plate.
Airpax 67L100 n/c thermal fuse?
bead with 2 wires 130K ntc thermistor
Setting a One-for-all so-called universal, URC 3445 in search mode ellicited
nothing in TV/VCR/Sat/DVD sub sets, as usual. What is not specifically
stated in the URC manual is after setting in search mode , why you press
power button, and then + to advance through the presets in the firmware, to
send power on/off codes.
I forgot to press Power in the DVD subset sequence and instead of
transmitting variously coded Power codes it must have sent various + codes.
Anyway on DVD code 1928 the "Keystone +" box popped up on screen. The "-" ,
Power , or no other button worked in this DVD subset mode. I intend getting
into the URC, remove ceramic resonator, and putting about 2V rms of sig gen
around the resonant frequency (done this before on a different URC) and see
if that opens up more buttons , set on 1928.
Failing anything more useful from that, monitor that "+" sequence on digital
storage scope should give the "50 KHz" pulse rate , coding clock rate and
combined model+code sequence for "+".
Then something like this circuitry of settable clock + CMOS switch settable
PISO shift register for model code (HCT165 ?) / settable delay SR (4557?) /
another settable PISO for buton code / gating a settable "50 KHz "
oscillator / IR diode / ( learnable remote for regular use ).
Changing switch settings for the button PISO from the known working "+" code
to the missing ones particularly for MENU and ENTER functions in particular
. Any flaws in this potential hack, as only thought experiment at the
ps , for the UK anyone know what is the cheapest souce of very basic
learnable remote controls?
fell at the first hurdle - the resonator in the one-for-all is 3 pin , 2
The default, not in vol or menu mode , for + is the keystone + function
The code , in the manual , for that via the RS232 link is
start code ascii 0 0 ,function r06 ,and end code (return)
30 30 72 30 36 0D
I wonder if any of that is reflected in the IR code for +
I've just captured that "+" code on a learning remote so I can play it at
will , rather than counting 105 steps on the URC . Whatever the code was the
URC sent is not stored in the URC as the same code.
So I've "learned" the 2 , to see what the difference is also.
Will have a look on a straight scope for the 2 basic timings and then DSO to
read the 0s and 1s.
I looked into the learner to see if its possible to change the resonator to
a much lower frequency to actually observe/count/time the code by swapping
IR LED to visible LED. But an 8 MHz 2 element 3 pin , I've previously done
that when I had no DSO access.
Even going down the LIRC route you still have to crack into the IR receive
hidden decoding to get started, if no listing. It would be nice to have a
general "all codes" sniffer , I was just lucky with this attempt this time .
Anyone know, if you had access to the uP of the device in question and could
put a nanoamp monitor or something in a supply line, whether you could zero
in , like cracking a combination lock
I thought I'd better try the learned code on the projector, but nothing. It
is all now making a bit of sense though. With the projector I was not
covering the IR LED with the other hand. With OFA 23 button push on its own
, isolated, produced no OSD. It now explains why I could only preset with
code 1004 and advance 23 steps as somewhere in between was a pre-code . By
elimination , by selective covering of the LED , it is step 6.
So programmed in 2 codes in the learner, "6" and "23". It now works with the
projector bringing up the Keystone + on OSD.
Codes are not sequential in the OFA , preset to 1004 and +6 pushes reads out
as code number 631 and 1004 + 23 reads out as 1928
So I thought set with learned "6" and run through all the OFA codes. If
there is a time-out reset between 6 and 23 it is greater than 2 minutes, as
I timed it . In TV and SAT codes something resets the 6 , in less than 2
minutes so something active in the wrong sense , but did not ellicit any
other OSD responses.
If the "23" preceeded the "6" in the OFA listing then I would not have seen
the keystone response , although (not specifically stated in the manual) you
can multi-step down as well as up.
Will set up scope and DSO tomorrow. I will also try learned 631 code and OFA
going down all its code banks to see if it hits something active before
those lurking reset codes.
Lools like a more general hack routine is to use 2 URC , preferably the same
model, for these 2 stage coded systems and automate the button pushes
I aalso powered-down the projector to reset any erroneous codes in the
system and then checked from fresh , that the learned 6 and 23 worked still
Looks as though the "23" precode of a Pioneer DVD is HEX
and the "6" code of a Samsung DVD is FF01
Perhaps a lot of the "23" one is ignored as junk and only the AA8 or 4AA8,
say , acts as precode.
So now to build a PISO with a bank of dip switches to code around the FF01
for the various functions (hopefully) and retain the learned code of the
"23" and use both to set up a workaday learner rather than fancy one with
LCD touch screen liability
Now where to find some of these basic Chunghop learning remotes
from a genuine co.uk rather than the co.hk or whatever, of that supplier.
"If something is too good to be true, then probably not true"
bought 2 of these excellent r/c, 1 for personal use to combine 5 remotes anfd the other
for cracking purposes
7dayshop Universal 6 in 1 Remote Control with Learning Function -
TV, DVD, VCR, SAT, DTT/TNT, CD/LD/HIFI/SKY/ SUMVISION MEDIA PLAYERS
Manufacturer reference: 5060207479424
Notes on that , uses EEPROM so your selection of codes stays permanently so
can be used via learn mode as an insurance policy to store, minus batteries ,
in a drawer for loss of main remote control.
To read out
long short short long blink means 001 code
0 is short blink
1 is long blink
Starts with a long blink
each non-0 sequence ends with a long blink.
(pressing 1,2 ,3 buttons is a much easier system to readout the code)
internally Q1 o/p LED driver
Q2 learning mode Rx
Will function with the 4MHz ceramic resonator removed and replaced with an RF generator
on low output level
For these hacks the Rx will operate correctly with the Tx code in the learner
but with "xtal" generator between 3.9M and 4.2MHz
IR remote control cracking
Returned to using Thurlby DSA524 for these non-repetative codes using the opto transistor receiver
described elsewhere ITF.
1 bit here is 10 pulses, wth 4 uS per bit and a 0bit is 0 for 4uS
The"6" code is
where a 0 is implied after each 1/2/3/4 etc block of 1s
6 is 6 ones and X is 10 ones
With 6in1 r/c audio / 056/Vol+
11521322X6165 same result as previous "23" code using the other learned "6" before it
Other Mitsubishi codes, while at it
where D is a long "double" 1 with no 0 space
27S4S117 , where S4S is space of 4 periods
126 / play
Not tried yet but to find other "6" codes
Try in search mode and follow each output with "23" code from the other remote
Made up an exerciser for determining minimm code , lopping off top or
tail of "6" and "23" codes and for changing the function code to feed into the learning remote.
Parallel to serial converter
Requirements 8 x 4021 , 8x 8way DIL sw, 4011 for inverter and
for gated clock pulse to output on the _CLK (without this gives long pulse, say 5 osc period pulse
rather than 5 repeated"1" s followed by a "0") , using a reasonably stable square audio oscillator
as clock source, fed through on the inverted waveform (_CLK), through the daisy chained
shift registers. 555 oscillator with 2x diodes and 2x presets for adjustable 30 to 50 KHz
pulses for IR LED via 51R to gnd, used p4 as gating input.
IT Rx from a vcr , L898, and 78L05 for checking as 5V pulse stream plus LED and
IR optotransitor checker for high f carrier stream
Push sw to go from parallel load to series shift.
The DIL sws for setting up the parallel data
eg 11222322412313211 up to 64 bits and the m/s arranged
to set the bit lenght and output say 3 "1"s for continuous 3 x1 duration output.
That gating the 40KHz oscillator to drive visible and IR LED
Working exerciser but not had time to delve into this project further at the moment
6in1 r/c uses simple 6MHz resonator so probably feedable from a sig gen to increase
range of codes
Uses 24LC32B (probably not replaceable with 128 as available memory will
perhaps stay at 32K but may be able to switch in another 32K if the switching is done when not active)
Casing clips on rear and retaining loops on the top at MACRO and between PIP+ and INFO
Exploring othe r r/c codes
There is one design aspect in out favour. To have many buttons/functions on
a zapper and not destroy ease/speed of use, the designer cannot add
repeats/checks & long code/decode sequences willy-nilly, as no one will put
up with a button-press response time of seconds for consumer use
Another thing I may get around to trying with this "universal"
is it uses a simple 6MHz ceramic resonator. Replace with sig gen and vary
the timings in search mose. I've done this before with another remote.
I have another one-time target machine a PVR that I could not crack the
coding of but managed
to get a replacement zapper for. If I can find a way of cracking that one
withoutout recourse to hindsight knowledge from reading the bought one, I
may (unlikely I know) be getting nearer a true universal system
I DSO'd 2 zappers
The cheap and nasty Ross sat TV one for
2230GKT-R 20.1 and Ross DVB-S 5010 set top box
that is easily corrupted by others and
easily corrupts other kit
for power button
for 1 button
so looks as though a lot of variability over the length (7x1 and 8x1 after
the 9) is something to do with non-uniqueness
The JVC PVR zapper that I could not crack via 3 or 4 "universal" is , oddly,
DVD power button
so seems to be just 111121 then permutations summing to 10 ,for the DVD side
These are for JVC kit on/off codes, I just DSO'd
1D71D1 (D is double width mark)
OneForAll 0133 code
Its worth trying other "universals" as they are not all the same , another
URC and 8 more JVC codes and 2 of a type of coding I've not seen before 1uS
single pulses and varied spacing.
To see why this ROSS R/C is so useless for intended purpose but may be
helpful in helping (perhaps shining ,gated ?, with a URC in search mode)
exercising "dead" kit, I DSO'd all the keys
So ident code of 91111111
1 / 136121111
up up 111191111
down down 115411121
other than sums to 17 (or else I probably made a mistake ) , 9 "blocks" and
ends with a 1, seems almost a random assignment
This zapper Ross DVB-S 5010 (rebadged Fortec ?) is probably a good start for
a hacking exerciser.
Pointing that 6in1 "universal" in search mode at the Rx , just for on/off
and channel change on the front display shows a response to at least 33
different codes assigned for different types of equipment. The inverse is
true but not so easy to put a figure to it.
Sony RMT D215P for RDR GX210
DVD power / 122421221111
OFA , more JVC codes for power on/off
324141(XII)5359 (XII roman 12)
0606 single short pulses and varying spaces
My parallel-serial IR RC exerciser/function code component enlarger, worked
, in that the waveform on the 1 to 5mS TB DSO looked and timed exactly the
same as the code I was emulating. Unfortunately I had forgotten that the
carrier circa 38KHz in not a square wave. Will have to a dd a couple of
steering dides to the oscillator and try again emulating a known system .
Reason I was not much good at electronics design, I always overlooked at
least one thing
The Logitech is only a URC with extra bells and whistles. It has the same
failing of all the others. If the maker does not have the code to laydown in
its library placed in ROM or via PC , and none of the other URCs have this
code to be copied across from. Then its as non-universal as the other
"universals", just costs more.
The only thing in out favour is the requirement for customer satisfaction to
have a speedy response to a button pusk.
And also the limited number of combinations possible if there are other
See my listing on the 10 th in this thread
There is very limited number of variants if you include a 111111111 subcode
because of the 9 blocks limit in this structure. A few more for 8x1 etc
That PVR I mentioned , where I obtained an original secondhand zapper.
There is a piezo-tick when the clock changes digit , or any othe rfunction
like choosing a letter for titling of DVD tracks etc.
Shining that bad RoSS zapper (clamped down FAV function which with hindsight
contained a large subsection of the PVR code) and the "6in1" "universal "
some codes elicited a few of thses clicks , unkown function if anything.
Reading out blink code fashion from the "6in1" one for something Alba 113
and one for something Philips 043
alternating codes D11DD11D1 and 1D1DD11D1
They both had the same use of double width pulses in the code. But now
having the genuine code for that PVR if must have been interpreting a double
width pulses as 11 and part of the ROSS
Hitachi 128 code, for ch 1
Gave up on the IR route and made a 3pin plus wide locator pin plug to go in at the
RS232. Worked fine getting into the menu via Win7 netbook and USB to RS232
converter under hyperterminal.
The image for red instead of vibrant was dull brown. Changing the settings brought
it to dull red and feint red and green , perhaps an improvement.
Reason for dull red not determined.
As though the red of the colour wheel has gone dark with age. Instead of
vibrant red it is dull red/brown. Improved a bit with up to the limits
changes to the colour matrix, now dull red rather than dark brown for a
saturated red input . Not a leads problem , computer in and then computer
out to a monitor is fine. Video source is similarly affected. No loose leads
inside. "Service" manual has no schematic of course.
Is there a generic cause to this as it seems quite common across different
makes and models?
I'm surprised these colour wheels hold together as they seem to be
individual glass sectors super-glued into a disc.
What an intriguing test.
Take one of the clear CDs that is used to
protect the bottom of a pack and, taking care not to get any
fingerprints on the disc, hold it so that the tip of your nose is poking
through the centre hole. Close one eye and move your head so that the
light from the lamp appears above the top edge of the disc.
I'll make sure no one can see me while doing it,
they may phone up the funny-farm.
(look, the loon has taken to playing CDs with his nose)
The lamp spillage looks normal white but
if such a spectrum test shows an absence of red it would be useful
I tried with some LEDs , the most convincing was "point source" non diffused
blue, just a slight bit of green in the spectrum with the blue. Even orange
lens over a neon was quite convincing. The more light the less convincing
presumably because the eye is overloaded with the most prominent allowing
the minor spectra components to be more prominent
Through one section of the fan louvres you can see part of the dichroic
mirror and that is distinctly red when the lamp is on , I wonder if the
dichroic plating? can break down over time and stop reflecting red and
absorbing it instead, ie passing red and infra-red light.
If I get inside again I may try mounting a piece of mirror in the lamp light
path and try reflecting a red LED laser into the light tunnel and
colourwheel and see if that brings up the red component/s of the image
Not just the bright light , but the cooling is disrupted by removing the
cover. I drape silicone sheet over wonen glass mat over the active area to
obviate shatter, intense light/UV, venting pitfalls.
The menu structure is that functions selected are designated by changing
blue to red lettering so are not that obvious as highlighted
The spill over light gives plenty of red on the "difraction grating". I see
this "blank CD/DVD" has a big clear patch on its surface, so those discs are
rejects from the first stage of production
Using slow USB to RS232 link .
To open Function Menu
Option menu. select language
send text for "enter" "+" "enter" triplet
to set function menu at header and then "enter"
Why is the red sector a denser colour than the blue or green sectors, or at
least on 2 makes I've looked at, or would more objective/scientific analysis
show the same colour density?
It seems to mean that if the lamp becomes under-powered or dirty integrator/
light tunnel etc then the reds become brown or plum , unless projected onto
a small screen area
For both makes, completely different sizes etc, the transmissivity , by eye,
of the blue sector matches the green sector .
This projector is used for graphical displays and red indistinguishable from
black at the usual diagonal is a problem. Cleaning the UV filter at the C/W
and cleaning the surfaces of the integrator tunnel has improved things a bit
but not enough. Orange or pink would probably be alright for this use
instead of red but dark brown or dark plum is not.
I just tried a scrapped dichroic C/W and it is quite easy to scrape lines in
the coating , say 10 percent , to pass something more like an orange colour.
Anyone ever done this to get round underpowered lamp problem and red/brown
issue? Would heat build up from the multiple scratches compared to just the
3 butt joins of the glass sectors , as made
Why dichroic for primary colours transmission and reflection of the
complement. In dichroic lamp reflectiors I can see a point in that it
reflects the visible light and transmits the heat through the glass to be
delt with by fans.
I cannot see that reflecting the complement is a by-product of reflecting IR
or UV back to the lamp
Instead of destroying NASA tech, I decided to open up the hole in the mask
between lamp and C/W, to see if getting more light will help, even if off
principal axis. At least I can put a mask back if there is too much colour
spill over from adjascent sectors, a more extreme version of adjusting the
C/W delay . That did not work because if red is brown then brown does not
become pinker by losing some red and gaining some blue from the sector on
that side or going orange by losing some red and gaining some green from
the other adjascent sector
C/W sequence B/R/Gn / Clear, 3 sc and 2 conn to release the C?W subhousing,
the ribbon is ZIF ,slide clip towards ribbon.
2 sc for mask cover, cut away metal leaving enough to hold the UV filter glass,
not much holds it in , as original.
Shine laser backewards through centre of projection lens to check it falls on the
centre of the lamp.
Integrator/light tunnel about 25mm deep. Cotton but too large , cut a piece of felt
and stuck as a U around a matchstick and then meths to clean off the duist.
If main, 3 wire fan stops then ps shuts down
Soon after switch on the red is corect colour but overall the picture at 4 foot diagonal is
dark. When full lamp comes on then the pure red disappears
To make a R/G/B scale test graphic. Set up lines in Graphpap.exe. Then in MS paint
fill in sections with R 255 and G/B 0 . Hue 0, Sat 240 and Lum 120 and grade down the
Lum to 0. Then in Irfan swap palettes R for G and R for B and copy and paste for
a triplet red-scale, blue scale and green-scale.
Traded off brightness for a bit better red with a rose-pink filter of 80 percent transmission.
Vaguely fitted over the projection lens with a large diam O ring, colour filter sample/swatch not really large enough but
image only passes through the central 70 percent or so of the lens for 6 foot diagonal projection.
Back to general IR zapper code monitoring
I've found references to their use as remote controls but now that LIRC does
not seem to be supported these days. Are phone cameras sensitive enough to
R/C IR in their own right ?
I've recently been playing with USB-RS232 converter and Termite (has 14400
Baud setting option ie 36KHz/10 x4 ) capturing of codes via IR receiver/
carrier stipper , works well enough for my purposes but if
there was an off-the-shelf phone-app that outputted the pulse train codes as
HEX test file or something, it might be more generally a simple and useful
The 4 digital cameras I tried all showed a flickering response to ordinary
IR remotes. When its dark tonight I'll try passing a transmitting IR over
one and see if it leaves an interpetable image over the 2 seconds of open
"shutter". As long as it/smart phones have the response time to pick up some
semblance of the typical 36KHz/10 of the coded bit length of 0.278mS rather
than the full 36 KHz pulses then there is a chance.
At the other end , the Hex coding would need gating with 36KHz into a
learner remote or something for normal use.
Well sort of worked. I didn't think of using the timer, at the time, and
restricted to about metre in front of lens and not being able to move the
R/C fast enough. Tasteful purple streaks. The one I used I'd previously
DSO'd and the first 6 .278mS bits of two 1s were clear , ie 100100 but the
remaining 90 bits blurred into 1 over about 1/25 of the frame width ,
partially because out of bad focus and perhaps overloading of the sensor.
Just for interest I will try again using timer mode and suspending the
camera so it will rotate about 4 revs per sec and hold the R/C about 3m away
and repeatedly pressing the button should catch enough part traces if not
one across the whole frame, raising and lowering the R/c also.
For timing and not just code sequence for anyone without a DSO perhaps feed
1KHz to another Tx LED mounted above the R/C in question. Whether a
smartphone camera would have short enough response time for a static use and
so a simple app is moot.
Could not get any useable image , returned to Maplin USB / RS232 converter
Added a small converter for the 5V output of an IR Rx from Sharp VC ... VCRs markings such as
6.B Japan L898
16G Japan L876 , with dimple collector lens , probable like Sharp 1SU60
Slightly modified the LIRC Thomas Breuer cct with 47K , 78L05, 4148, 2x BC184, BC214
to convert /buffer 5V Rx output to c +/-6.5V of the RS232 unit
25 way , D, p2,p3,p4,
p3 data, p2 "0v", p4 (+v)
If using grounded scope disconnect the laptop from its ps , with USB-RS232 adaptor in use
Downloaded Termite "Hyperterminal" as gaives option of 3600 Baud (for 36KHz/10)
Use Terminal Font if Hyperterminal
Use XVI32 to convert ASCII (Terminal font) to HEX and block copy via Clipboard to TXT file
Termite you can select HEX + TXT, another advantage
RTS/CTS/ , None, Hex View no word wrap.
then convert HEX to binary and gate 3.6KHz to return to IR Tx code
Needs decoding for 2400 Baud as wrong timing, 3600 Baud not working. Added another inverter
to the circuit to give a negative output for converting to HEX , but ran out of time playing with this/
working out how to use 3600 baud setting
Converting a 200W discharge lamp video projector to LED
Assuming over-riding the opto couplers to falsely confirm to the system micro that the arc is struck and the lamp is lit (maybe requiring a delay) firstly, with the lamp ps disconnected.
Then I intend in the first instance to buy 5x 1.2W white LEDs, 3500K, 20 degree, for proof of concept. Assuming that sort of works then get perhaps 10 more, going down to 2700K or 3000K or perhaps (unlikely) 4000K and lenses to colimate to 2 degrees. Set inside a reversed conical silvered glass of an ex-lamp to direct spillover light into the colourwheel/light tunnel aperture.
The intended LEDs are 11x10mm footprint so can be mounted quite close to the colourwheel(for 5 anyway) on a spherical back mount. I may as well retain the original fans, perhapps knocked back a bit for less noise intrusion later on.
When coming to scaling up I originally was thinking of using a sectored curvi-linear silvered reflector from PIR units (with faned air cooling) but have since come across 2 degree lens converters for these LEDs so may as well go with them and shine directly from a larger spherical backing mount, directly to the colourwheel aperture.
I'd be interested in any suggestions or comments other than of the I would not bother type of replies. Anyone happen to know what the light wastage proportion is of a non-ideal paraboloid reflector and non-point source discharge lamp is? I'm aware proper LED projectors have active drives to RGB LEDs and not colour wheels but there are a lot of ink-jet-printer-syndrome surplus HD video projectors around with too expensive short-arc lamp costs to replace
Some bods been here before with converting a couple of types of discharge lamp converters
If anyone is interested, my exploration of inside a standard domestic GU10 LED lamp (to see if they were all in
series or mixed series/parallel)
Breaking in:- hold the bulb in a glove and heat the dome cover with "low" temp hot air and prize off with a needle. With old soldering iron destroy the epoxy join between the , not obvious as silvered, pcb to the lamp housing. The slight greeen colouration is due to the reflection of the green dye of the pcb which is not glass fibre reinforced it seems, maybe epoxy substrate only. Prize the pcb away.
No glass breakage at any stage.
Overlay of this one marked JH-GU10-20
HV ac side 1M//0.33uF 400V dropper and 1/4W resistor size fuse/inductor/fuseable resistor? pink colour with red black brown, or reverse order, colour bands, about 0.4R to small SMD MB6S bridge rectifier.
LV quasi-DC side SMD 510R dropper to 20 LEDs in series.
White ceramic cap is cemented to the glass of the lamp.
Bench ps 50V across LED string and 510R all LEDs just lit
54V and 0.5V over 510R and about 2.6V over each LED some sort of low level brightness.
With 75% mains (240V that is) 6.5V DVM dc over 510R
or 5.5V DVM ac over 510R
100% mains 9.1V "DC" or 7.2V as "AC" reading over the 510R
Back to projector
Its not possible to do any calculation as to brightness because it is impossible to find the true loss of light from a conventional distributed ie not point source discharge lamp source and non ideal reflector, I would guess that 2/3 of the rated light does not get into the apaature and then how to calculat the proportion that is at such an angle to the light tunnel to the active chip and multiple reflections that little of that gets to where it is wanted, axially along the 2 inches or so of narrow diameter light pipe .
As for colour rendition it is unlikely to be any worse than the current situation of having to place a rose-pink filter over the projector lens to get some red into the image as the lamp must be too far into the blue end of the spectrum to be compensated for in the setable timing of the colourwheel.
So in both cases very much a suck it and see, seat-of-ones pants situation, but worth a go, if you've seen the price of these supposed replacement discharge lamps
I've had another look at the product data and the 2 degree lenses are 4 degree , in normal terminology
I don't know what chromaticity means but for the 3500 deg K version a Cx of about .4 and Cy of about .39
A bit more graphic the spectrum is continuous and smooth "bell curve" peak shifted 40nm from 550nm of the standard eye response curve to 590nm
and the 50% points broader apart at 150nm compared to 100nm of the eye and a 50% down peak at 460nm which I suppose is the potential bugbear for such a lamp conversion
LEDs, I did not find a uk/us URL
All the discharge lamps in video projectors , I've seen have an electrode and squashed , not optically conductive glass, axial exactly in line to where you want the light to go.
With directional 20 degree central LEDs, and lensed 4 degree ones off-axis. will direct most of the energy directly to the half inch aperature , without any reflectors . Reflectors just to mop up spill over .
I can see some high power red LEDs being added to counter the blue sub-peak
I've not found the data out there but I suspect the vast majority of the light available to a gobo setup just does not get into a light tunnel setup , so if you can direct all your LED lamps into that tunnel then the overall requirement is much lower than normal ratings would suggest.
The use of this video projector is for text and graphics so colour rendition of the likes of flesh tones is not too critical, very rarely showing any video as such.
I've moved away from thinking about using lenses as they restrict the footprint size. If I double stack the LEDs then a spacing of centres about 8.5mm is possible and so 14 LEDs in a 32mm diameter. So can be quite close to the light tunnel and 20 degree beam spread is fine and outer LEDs approach angle will still only be 30 degrees or so, and a reasonable proportion of that will reflect only a few times and substantially get to the DLP chip. Will still try 5 LEDs initially.
Will power up individually at only 100mA or so and set the angles of each individual LED for maximum brightness at a simulated window with a photodiode , then wire all in series and make a more substantial back mount before transfering to the projector
The lamps used in video projectors are ,in effect, not compact. The direct light path from the reasonably compact source is blocked by an electrode and non optical structural glass lump, relying on the mirror surface of the light tunnel / light pipe to average out/balance-up the light coming in at all sorts of angles, off the parabaloid reflector
That Osram with built in lens is discontinued, only came out 3 years ago.
I'm getting 7 of the LCW W5SM, white 2700K 120 degree and matching 7 off 6 degree hex shaped lenses.
The off the shelf 7 cell hex cluster lens is for parallel , not focused.
Using a 5 to 6 inch ball as a jig I'll combine the 7 as a focused cell.
With the lenses at 85% transmission, brings the 75 lumen per LED down to 64 lumen, so x 7 =450 lumen. Some heatsinky type protrusions added to the LEDs to catch the fanned air.
When new the projector was rated 2000 lumen, although still taking 200W , the light output is now much less. Combined with the poor optics of these lamp setups I suspect fully directed into the light pipe 450 lm is not much different to the present discharge lamp situation. Still plenty of room to add another 6 plus lenses around the periphery to bring up to 900 lm.
Playing around with a scrapped colour wheel dicroic disc and assorted white LEDs the red transmission , to eye anyway , is a lot lower than G and B.
So I will get a number of red 100mA 5mm , 15 degree standard size LEDs to add red, perhaps 6 at the interstices of the lenses, mounted to the rear and between the SMD LEDs plus maybe another 6 around the periphery to infill.
Good progress with the silvered "cone" for mop-up, front of PAR lamp removed and bulb removed, about 3/4 way through grinding through the thick glass with cintrided disc, to remove the barrel part. Previous attempt with thinner glass photoflood failed. If I was brave or had a load of these sort of lamps, I would try the old bottle cutter routine, freezer spray and a ring of "fuse" wire around , fired up for the cut
Well that was very successful, a job I've never done before. Not a cone but the nearest I could find with
a good silvered internal surface and right sort of dimensions. The thinnest part of the glass 4.5mm thickening to 7mm as 12 flutes around the stem,
but a neat ground-glass cut. Cut a continuous circle around about 2/3 in then concentrate on a pece
about 15mm long until breakthough and then progress along each side until not enough glass left and the
remaining will crack through
So I have a spillover reflector with 70mm internal diameter available for the LED assembly, down to 21mm diameter opening for the colour wheel aperature (from memory about 12mm) and 43mm axially.
It will be a couple of weeks before I can get some time to convert the projector . But in the meantime anyone any ideas how to measure the intensity of the LED assembly at the axis and focus of the array, with any sort of accuracy (no known "standard candles" etc in my possession)?, for anyone else coming down the same path. My uncalibrated luxmeter , I doubt goes that high plus probably would melt, as would any of the plastic neutral density filter I have and placed in the path of that beam
Its more bodging than optics. I'll take a few pics along the way.
I've just noticed on the LED lens makers datasheet if you put 7 of their 1.2W white light LEDs in a specialised version of one of their lens structures , as I will be doing in a different manner, and directing to a 12mm waist , you end up with 3,000,000 lux there . All I need then is some deuterium pellets.
I think I'll need more than neutral density 4 stop Lee Filter , I just tried black silicone rubber sheet and that is more like x0.001 light attenuation with an odd colour cast from white LED
has some cell cluster focusing devices.
One of those that if you close focus , you can start your campfire without having to rub two boyscouts together?
an example of the focus mechanics shown in here
I now physically have the LEDs and lenses. 2 immediate problems, both active faces of the lenses are dead flat. In the pdfs it looked as though the front faces were surrounded in a ring and I assumed (pdfs not clear pics) the 4 corner holes of the LEDs would mesh with pips on the rear of the lens.
They are made for mounting to pcbs not the other way round, so no provisions for that. So I have to find some 12mm or so diameter thinn rings to align the front faces to my 5 inch cistern ball valve float as spherical former, and make some sort of jig for aligning the LEds to the lenses , plus fixing them together
I'm not used to optically pure plastic, I would have sworn that examining the LED side of 2 of the lenses then it was flat faced, but not touched the surfaces, and they contain recesses.
I now have some rings for placement.
The datasheet does not explain the normal placement. LED soldered to pcb, a non-optical plastic holder placed over the LED, quite loosely. The holder is .4mm undersized so forcing the lens into it, then compresses the other end around the LED with little pip under each corner, but no use made of the 4 holes in the corners.
I think I've worked out how to adapt those holders for my purposes.
Supplied in small numbers , the lenses are supplied pushed into the mounts and again they don't say how to remove them without damage to the lens.
First one bodged by placing over a couple of metal plates providing the across-flats 12.9mm gap to then push down as detailed below. As stated previously you lock the holder to the pcb in normal use, by pushing the lens into the holder , so first they have to be separate.
Then with 1 free holder, place back-to-back castellation to castellation over the next one. Place a flat piece of rubber into the slot part that takes the LED (so not to damage the thin lens wall) and push down quite hard with a rod, until the lens pops the 1mm into the precise recess of the other holder. Remove the second holder by hand and then grip the lens across the pre-existing mold marks and pull out from the first holder with serrated edge pliers.
Well that is my method for Osram Golden Dragon lenses made by Polymer Optics Ltd as supplied by RS
The focused 7 cell honeycomb lens came together well. Doing the maths and having a wedge tapering to .3mm seemed ok but trying to adapt plastic to those sorts of dimensions is not practical. Having to mould my own mounts.
The 2700K LEDs are noticeably cream yellow colour to the eye and seem to have supressed the blue peak, too much?
Relative intensity through a dicroic colourwheel
even the reflection off the blue filter, ie complement colour of yellow , is noticeably brighter than the reflections off the R and G sections.
I was expecting to add 5mm red LEDs but looks as though it will have to be blue ones. Won't know for sure until the video projector has a reasonable block of time to get inside to mess about.
As my mouldings to fit LED to lens are egg-cup shapes , a 3 part mould required, letting epoxy harden before running off 7 of them, so far so good.
I was impressed with just trying one LED at 1/4 power of 90mA (no heatsinks yet) and 5 feet away , bright enough to read by just.
Add a lens at the correct position and a neat bright foot by foot "pixel" , of the chip thrown on a screen 5 foot away, probably as bright as the projector showing white (DLP losses unknown).
So 7 of them and 4 times as bright would illuminate a 5 x 4 foot screen , with that brightness. If 50% loss in hte projector then add another ring of 6 , room in the "funnel" reflector for them if required.
I wonder what a gobo type thing made of a matrix of single RGB LEDs , with lens arranged to throw foot x foot squares with 6 inch overlap , would look like if driven with graphical/animated "video" pulses
( Sun data 0.0079 W/m2 per Lux , 1 Lumen= 1 lux at 1metre)
Converting discharge lamp to multi-LED for video projector (maybe)
Story with pic so far, adding to previous thread text a couple of weeks back
7 lenses , each placed over a 12mm brass cup washer, to seat over the spherical surface of the 5 inch orange ball valve and flat lens fronts (set in a wheelbarrow innertube of all things)
Tape to quasi-hold in place and glue with superglue spots when posistioned
One of the lenses with supplied white holder on left and 2 of my abandoned attempts to convert the holders to rear mounts for the LEDs, mm graph squares
Pic of PAR lamp innards , ES base on right with small fuse inside
Pic at 90mA (140 lux with black silicone, 6400 lux with 9 stops
KISS, pipe cutter to cut down a large blue barrel pen to intended even lengths then adjusted by countersink tool to fit over the coni-spherical rear of the lens , for proper length and provide a temporary swivel for each , to active focus each before setting as the 7cell. A basic pcb copper 20mm fuse holder soldered to each LED as a heatsink, then LED set centrally over the
blue and hotmelted in place with yellow hotmelt. Wire and solder all 7 together.
I did not think about sliding on some sleeving so each link wire stripped back a bit in the middle for croc clip connection. Set some perspex correct height over some paper with a square marked on it and clamped down . Set first LED and mount , held roughly in place by a light clamp. Then powered up central LED at low current and adjusted mount for square die image to be central to the paper square, hotmelted with green hotmelt into place. Powered up next LED only and manipulated orientation and setting for its image to lie in the square also . Repeat in turn for all of them .
The luxmeter I have is not to be relied upon as far as calibration. With some neutral density filters of 4+3+2 stops = 9stops or /2^9 luxmeter read 6400 lux at the centre, this is in non-linear section of the meter above 3000 but this same meter read 2,100 lux with same 9 stop filters, horizontal outdoors when insolation from the sun on a cloudless day was 860W/m2 (includes UV &IR from nearish weather station). Non neutral density black silicone rubber over the sensor read 140 lux, the same as with horizontal and the sun (perpendicular to the sun it read 300 lux that day). This was 90mA , in final use with fanned air movement would by 4 times that, so about 6 times (plus non-linear correction , not yet determined) the horizontal lux level of the sun that day. Still room for 2 rings of 7 each could be added spherically around the 7cell and chromaticity varied if necessary. I've now got access to the video projector for a try out , only one opto for "good" signal out, the other 2 must be for lamp trigger , then on receipt of lamp good then lamp full on control applied.
Assuming simple on/off opto use , have to check not PCM or something awkward
repeat image with added stopping of camera lens , trying to show that the 50% brightness contour image lies within the square
repeat with only 5mA of power to the LEDs difficult to photograph as available light only. Note the appearance of some graph paper markings inside the corners of the square
Bad news for the Jeremiahs and Thomases.
I have a converted HD video projector , as far as I've gone so far. Hacked the control system so the original lamp supply is disabled and the control is fooled into believing the missing discharge lamp is lit.
Jury rigged at the moment so the fans not in the right place so running 7 cell at only 75mA. But with a colour bar signal a good image to 2.5 foot diagonal in a darkened room. Good colour gradation across the luminance scale for red and green , the blue the last 3/10 to black are black but that was to be expected with 7 yellow-white LEDs with low blue component and existing adjusted projector settings trying to subdue the over-prominent blue and lack of red component.
With fans back in position I may try the full x5 350mA rating
I well set the projector back to factory settings via rs232 and power up again before settling on which flavour of white LEDs to buy this week.
Probably 12 more for the next ring , mid-white LEDs for 19cell and 14 times the power for the normal 5 foot diagonal.
There is room for another ring beyond , 20 ?, but worsening angle to what TI calls the integrator rod but I call the light tunnel and significant expense, but even then a lot less that 400+ squid for a genuine lamp or 300+ squid for grey-market one
This arrangement has much more latitude as to precise alignment in front of the light pipe than the parabolic original. I assume because no direct axial light was going down the pipe with original, instead a tight focus at the pipe entrance with a high angle of incidence for the majority of the incident light.
In the process of getting 12 more 4000K, this time, 1.2W LEDs and 8 deg lenses plus some directional blue if I need them.
Also sorting out a compact power supply, converting a laptop or similar SMPS
Some pics of the 19cell concentrator compound lens and LEDs
Some pics of the 19cell concentrator compound lens and LEDs
and very low , few mA, for camera capture
to the eye the central 7 , 2700K ones seem even more yellow than the 12 x 4000K ones
At 50mA per cell the RGB balance has come up to
R 360 lux, G 450, B 230 through a colour wheel filter
Today I will add a ring of 12x 15 degree high current 5mm blue diodes around the perimeter to bring up the blue (if necessary) before I mount in the projector . Then baffle off to divert the fan air over the LED cluster before trying at 350mA. From further reading of the data sheets, as fanned, could be pushed to to 900mA with lamp life derating , but I'd mount a 1N4148 as a temperature sensor in the clust before going there . Would only be used at the high rating in a couple of summer months , I hope. Have an adaptable 42V 2 amp compact SMPS that I'll add some polyswitches in line with the 3 lines of LEDs and droppers . May have to move to constant current , dependent on monitoring the strings at high current
This 1080 projector was 7 GBP at a radio rally because the replacement lamp cost 450 GBP, the ink-jet printer business again.
The logical way of making a 19cell cluster is to mould the compound lens in one. The geometry is not complicated I think, just the moulding technique and optically clear epoxy . The recessed lens part of each lens looks like the dome of a clear 5mm LED and that is set in a non-tapered cylinder that collects the light, perhaps via total internal reflection off the quasi-spherical rear outer surface.
If good geometry/moulding then perhaps no need to individually focus each LED.
The main problem is converting the projector, finding the opto-isolators and defeating their normal control function
final run of the Mark1
the original colour bar chart, sort of chrominance on left of white stripe and luminance to the right
RGB approx values
plus odd stripes etc of white
slightly skew and not a proper screen , off white cloth as a screen,
with 1 metre ruler slung across but as dark wood not obvious as a ruler and my camera does not like low light values. The green to blue and red to blue in the lower right corner is an artefact of the camera unless brain/eye is wonky too.
Just as well the human eye can accommodate low light values better than my camera
projecting a white image, the received intensity is the same as a regular CFL lamp marked 850 lumens at a distance of 5.2 metres.
Not really bright enough so I know what the mark2 version will be.
The problem is getting the light focused into the 4x6.5 mm tiny aperature of the light tunnel. Those hexagonal lenses are too large, so by the time you have 19 the off-axis angle to the light tunnel is getting too great. Will visit a bead shop and try to find mirror glass beads with about 5mm core that will accept sliced and polished domes off any old water clear LEDs as lenses set in the mirror tubes and mounted over the 1.2W LEDs, and individually align to the aperature via an intitially tiny amount of hotmelt string to the LED rear and then a matrix of lacing cord and epoxy attached to the 20mm fuse clips added as heatsink to the rear of each LED. Need as much unobstructed air flow through those vanes as possible.
Defeating opto controls to and from the Mits ps
p1 -ve HV
p10 control on/off , high for on via 27K
p8 raw 18Vsupply to 18V via airpax
p9 18V regulated
2 thermistor 7R NTC
red of red/brown to HV lamp ps is floating 0V to both heatsinks
discharge large C907 before handling
lamp ps ST UC3842B control via 47R and 2 SMD Q to gate of MIP0221SY also uses uPC324G and uPC842
Disconnect HV feed to lamp ps . Press power on, 2 x green LED as normal
then red and flashing green and no colour wheel, then no fans
Fitted 1N4148 to replace opto coupler, permanently connected ok, no need to switch it in/out
Added 2x LEDs +470R and the 1N4148 to the control line to rest of the machine to defeat control to/from the
Cloth tape over base under the LED lamp to avoid any heatsinks touching chassis
light tunnel internal measurements 6.5m x 4mm
Initially made a flatted cone from tinned brass, then polished , unfortunately
ended up 45 degree sides so double reflected . Redid with shallower angle then
of course compromise of cutting off light v funneling light with too many reflectionsRemote control for the XD200U. East London Components RMC 13290, uses SONIX SN8PC20SG,
correct coding for most of the functions. Plectrum at middle
of long edge at first.
This one , I had to break into for curiosity, uses a SONIX SN8PC20SG full data out there, and little else. Programming via 4 pads accessible via the battery compartment, so 4 data lines for that and the other 13 data lines for presumably 5x8 matrix 40 keypad. What is the process of going from LIRC or some other R/C database or OEM or learner r/c from an original to getting data inside one of these one-time learner R/Cs. Or some other more general process?
A bit more
"rm series" "remote control" in google images shows there is a generic physical form of 40 keys or higher keypad counts and within them subsets with differing legends for DVD/TV etc , for hundreds of different makes and models within those subsets. Just the data must be different inside and a different RMC number. Not all functions are necessarily present eg my one has no PiP function that is on the original r/c but never used anyway, but it does have the vital Menu and Enter key.
What an enterprising idea on someone's part. I just wonder what lies behind the data accumulation/distribution, part of the dark web?
If they pulled the data off "universal " r/c s then it was not on one of the 5 or 6 "universals" that I tried.
A market stall near me has a few boxes of hundreds of assorted used r/cs for 1 or 2 quid each. Perhaps something like that or a large charity shop that gets hundreds of donated units with remotes , but are not sellable, is the source of the data. But usually like that only the maker name is there , not the unit model number that those r/cs relate to
Panasonic WV51E CCTV camera
No o/p and also loading the monitor supply.
Small dropper to the heater had touched the chassis
shorting the 9V rail and also knocking out the
470uF,35V coupler / de-coupler at the output.Note
the camera needs a frame sync source (to sync all 3 cameras) from the monitor /
supply as well as 12V
Pyle PLVW 1342R
had been dropped
12 inch diagonal called 13.4 inches, 12V dc supply as used here but may originally have had a
CVBS video input
made 2005 to 2009
Build into an expensive bit of specialised commercial kit and the original
outer case of the display absent.
Main pcb overlay marked "Sharp8 pcb"
main chippery Techwell TW9903
3x AMS1117 3.3
SMD diodes, 2 marked
xtals 12M, 7M, 14.3181
red 12V, Lt Blu 0V
yell , Gn + Dk Blu
Will flash up pink backlight for half a second but no LED
If no gnd continuity to Al foil gnd at inverter and gnd conn at push button
bank to steel frame then LED will not light and no sustained backlight.
Normal action is RED comes on at powerup or set to standby
and blue LEDs in use.
Sahara DP3601 video projector,2006
Bright orange, 2 of these , minus the security necessary top control panel
for each, not found a pic of them yet.
15 way connector for a panel , that when removed, disables the projector.
Any chance of hacking them, there is a IR receiver lens but that presumably
would be disabled with an absent top panel? presumably most of those 15 ways
will be a diode matrix for control buttons and anunciator LEDs.
Little info found on these projectors but Google images suggests Delta and
similar model numbers maybe a related badge name. Googling the lamp numbers
has not shown up any model Xref
One lamp is physically blown, the electrode , external to the axial quartz
glass stem but in the central bright quasi-focused beam area is broken, not
actually where it emerges from the stem. Surrounding the break , and also
the supposedly working one, in that area, is a neat looking aggregation of
spikey white crystals around the conductor. I've not found an image of a new
bulb yet , but could there have been a piece of silicone there that has
degraded to silica crystals (asbestos-like)? perhaps creating a hotspot for
the lamp current and then failure, actual and potential, at that point?
Any pointers to a procedure for hacking the control-defeat function of these
a pic of the projector and top control panel
"Each projector has a unique serial coded panel and it recognises only that
one; replace it with an unauthorised panel and the projector will not work.
Therefore, should a thief obtain even one panel, they will not be able to
re-boot any machine but its manufactured partner."
so presumably the only chance is to find some powerFET control switch or
something simple like that
Comes apart quite easily . Seems to have 2 separate smps , one for the
electronics proper and one for the lamp only.
Of the 15 lines 11 go to 3K3 pullup Rs , presumably for key matrix.
Of the remaining 4 , one is gnd, one is clock line of an I2C 24C08, one to
its data line and one to its Vcc
So what sort of thing goes in the missing control panel, like a car radio
removable control panel, linked to those 4 lines?
Discharge lamp of 2000 Lumens from video projector
a new one
break at the crystalline growth
pictures showing the electrode that is axial to the glass stem , all
contained within a parabaloid reflector so at the central area in the area
of heat and light that becomes the focus in the optics. The discharge lamp
proper is farther along this glass stem , outside of the pics , difficult to
show as fully inside the parabaloid down there.
The needle-like crystals do not show up in my pic unfortunately. White with
a slight tinge of green. Failure is at the mechanical stress point of the
bend and I assume the crystal build up collects heat as well to further
The lamp cooling air flow from fans passes over this crystal area. The
diamond pattern is a double spot weld to the metal that then connects to the
supply, presuambly a diferent metal. There is a dusting of this white
substance , not needle crystals,on the electrode down to this diamond but
not on the continuation wire.
What is the chemical of the crystals and the mechanism at work here?
on the outside free-air side, perhaps I used the word electrode wrongly,
I assumed it is tungsten wire to take the heat, ie not copper. If Y cleaned
off the crystals and remade the break in the wire then the lamp would work
as normal. Another lamp from a similar machine is still in use , but that
too has this crystal growth and presumably will fail at the same point , at
some point in the future.
Perhaps I used the word electrode wrongly, I assumed it is tungsten wire to
take the heat, ie not copper. If I cleaned off the crystals and remade the
break in the wire then the lamp would work as normal. Another lamp from a
similar machine is still in use , but that too has this crystal growth and
presumably will fail at the same point , at some point in the future. On
close inspection of the glass of the pillar, the surface is not smooth
blown/melted glass but a bit faceted as though small parts have fractured
off, perhaps its that which has transformed into needle crystals
Before using the working one, also with this crystal growth, I'll do that ,
and report back what sort of surface to the tungsten? wire at the bend and
the state of the glass at the end of the pillar. I assume its quartz glass ,
to take the temperature. The clean image of a new one is taken from a
manufacturer's www site, I only have 2 , one pictured with broken wire,
both with crystal growth
The problem is not in the active bulb part of the lamp, just where , I'll
call it a conductor wire not electrode, emerges from the glass stem that
contains this conductor that then goes down to the lamp proper.
The break, with the crystals around it, is external to the evacuated / Hg
vapour active section of the lamp. Spot weld or crimp over the break and the
lamp should continue working for another 1000 hours or so
These lamps need a trigger/ignitor voltage of something like 3 to 5KV, once
the arc is struck then something like 70V,2.5 amp DC maintains the arc. Must
have been building up over a few hundred hours of use.
Scraping back the surface of both types of conductor, neither is copper
colour, normal metalic silver colour, I would not know how to tell if
tungsten or kovar or whatever.
I can now look at the break area and glass pillar under a x30 microscope,
after soft scraping off the crystals. There is a neat conical end to the
glass pillar, with the emergent conductor at its centre. The surface of the
cone looks very neat as though etched , and not part of a production molten
glass process. I suspect microscopic thermal changes at heating or cooling
between conductor and glass cause micro-shattering into slivers that
somehow, unknown mechanism, stick to the conductor, rather than get carried
away in the fan-produced cooling airstream that passes through this area.
Over time and use this fracturing process grows deeper and forms a cone.
Perhaps they are glass slivers rather than crystals.
Where the conductor break is, the metal has belled out before breaking
Curiously the place with absolutely no glass slivers is inside the cone.
The slivers showed no effect with magnet , nor the conductor with the break
in it. But the conductor beyond the diamond double spot weld, ie away from
the most intense light and heat is attracted to a magnet.
Humidity unknown , other than used and stored in a heated habitable brick
built building, unlikely any condensation
Although no schema for this unit , googling on the 2 main
"video " processor chips and filetype:pdf threw up a few detailed
schematics, to then targetted probe under the big BGA, to check traces.
Is there a special probe for doing this, thin , tough, insulated , with
graduations for indicating which row/column depth from the edge ?
I have some .08mm tungsten wire that is just about stiff enough as a probe ,
for this Ti DDP2000 BGA could have been larger diameter, laquered over and
just the end ground down clean. A refinement I suppose would be contrasted
markings along the laquer for row/column spacings, I think up to 5 deep in
some areas of these
DAXX DPP537 Plotech
3 sc under
Lift grey rubber at lamp lid
2 sc at lamp housing to remove lamp
remaining 6 sc and lift 2 part middle top hinged section away, 2 sc under, top
orange comes away,release ribbon
Bare lamp Osram
P-VIP 180-230/1.0 E20.5
#2 fans kick over for 1 sec
ST LD18 U800? , 0,1.8V,3.3V
3 sc for front panel
2 sc inside base of ps, 1 small sc under rear panel
main pcb sc , lift main pcb enough to get inside ps and release the DC conn and angle enough
for ps fan to clear the main pcb
4K7 to Vcc
HDCP keys for encrypted video
1&2 yel & red, !V DVM-D either way
1&2 Bn &Y
P12NM50FP double R logo for Samsung
SEC 546 TCS 8
WKJ050YN overlay U500 , 7x9 DRAM
Ti TVP5146PFPNTSC/PAL decoder
Ti 5AD5KiK CDCR83
16p PIV PI5V 3300E Z0532
AU230 Quad OR
20p M6239FP DAC
con306 4w to/from ?ps
16p AD1034A RQZ, ADM1034 fan control
24p AD 8183A #0541 761307.1 MUX
05DZ11,p3 via 10K
MX 29LV160CBTC Flash memory?
ps 2 sc fixing
cut upper Cu tape , and tear lower tape
con306 to mains ps for small case-mounted thermal monitor SDA and CLK of 1034A, also
connected to top panel and internal I2C
6p SMD T100 thermal monitor, Ti TMP100 I2C SOT23-6
p2 to blue, p1 to red, p6 to grey, p4 to white
C between p2 and p4
DDP2000 BGA connections
C22 to TP702lamp lit
B22 TP703 lamp on
C17 power on
C22 to CN601 via 110R, R645 marked 05A (ribbon con p8 to near C19) near to B19
Front IR-1 C29 , C27
lamp uSW to CN600
thermal fuse? under the lamp
bridge contacts to trace back to B19 case-det
out power-on C17 ribbon p4 via 40R
U601(5)SDA to A27 SDA-ch1
U601(6) , SCL0 , B27 SCL-ch1
B16 power-on via keypad
CN605/4 via R648 to U602 between Vid ins and outs, CCK topcode
back to "CEK" U607
The nissing control unit probaby has the 9 keys as a 3x3 matrix.
Probing under the large BGA with tungsten wire , so probably errors of placement, and using LG RD JT91 service manual
as a guide, for the 15 pin conn to hte control unit
p1 DDP2000 B16, KEY IN 0
p2 C16, KEY IN 1
p3 B17, KEY out1
4,A17,key out 0
5,C17,key out 2
unplaced D16 key in 2
p8,B19 fan tach2 in
p9, BorC19, fan select out
p10, A19,480I/480P ip
p14, p6 of I2C EEPROM
p15, p5 of EEPROM
Unplacedconfused pinnings probably relate to 2 LEDs , one or both
bi-colour perhaps/ different supply or return line?
Lamp repair, 2 the same. Cut away the original wire at the
fixing tag on the reflector.
Some heater wire from a 200W, 110V heater, 0.4mm diameter
st/st nicad strip curled around doubled up pair of .4mm .
Pull out the "former" and transfer to abraided end of the wire
stub at the lamp end. Crimp with cut-down by grinding ,pair of small mole-grips. Pull wire through hole in the reflector, wrap around
the tag a few times and crimp a bit of copper crimp from plastic
removed blade crimp.
Samsung PS-50C96HD Plasma TV
First large "flat" TV I ever look at and its a plasma one. Dust never
cleaned off the rear grills and a visibly blown cap in the ps. Anything to
be aware of, concerning any likely stored voltages/ safe discharge/
reconnecting headers as there are loads of 275V rated caps on the other
boards, before/after removing the ps .
Other than never hold / lean against the front panel, and no bending of my
back for lifting /carrying, any other service advice specific to large
Cycles continuously through coming on and relay clicking off , starting up
again,shows on screen when active
CHECK SOUND CABLE
Dates of 2007 around it. 2 swollen/burst + one poor ESR all 105 deg C ,Low
V, replaced, not had time to power up again. At least there is plenty of
space in these sort of things for different footprint or taller caps as I
like to replace with next higher voltage.
All replaced 3 CX402,403,411 were Samwha WB , brown and silver , brown maybe slightly
31x49 inches , 40Kg
Came to fit ps back in and although an extra reinforcement ground solder tag
, no pillar under , so no missing fixing screw, as no chassis pillar
mounting point. Working order again
Sony EV S700
4 allen sc to remove front cassette flap, 2sc under front panel, 2 sc either side of top metal brace, mark and undo
CN20 , 4 asymetrically
disposed sc to remove cassette carrier, unclip the end of tape sensor at the 2 inside plastic clips.
Depress nib inside the cassette tape to retract tape as VHS.
At leaf switch push white catch/latch towards the outside
Broken cog stopping tape retraction although cog is on the take up side. Remove brake arm with a finger
ove rthe circlip to retain while prizing off. Lever off the small closure to
the spindle, plastic broke , tried fashioning a snap reivet but failed , original was spherical so in effect wedged
which a snap rivet stem is not, so used a dot of hot melt on reassembly.
Spool assembly interference fit of the 2 main parts , use parallel jaws to grip and twist
Reminds me of a notorious generic fault with Philips car cassette players of years gone by. A steel disc set in a hard plastic periphery, differential expansion with cold or heat (never discovered which with the Philips). In that case periphery was of pulley form taking a drive belt, this time cog form mating with a drive cog.
Part of the take-up spool structure of Sony EV S700 video8 VCR from 1985 , used in 6 channel audio mode, so again a functioning m/c required to salvage audio from a tape archive (like the DAT m/c in recent thread here}.
I placed the broken away quarter of the ring gear back with the remaining 3/4 ring and there is a stress-relieved gap of about half a tooth. Any advice?
At the moment I intend grinding back a sliver from the steel, along its exposed quarter, to allow the bits of plastic gear to align again, and allow non-jamming tooth engagement. There is a bit of a ledge to the plastic cross-section to allow for some glue along the length as well as the broken ends. This spool drive assembly contains a magnet that "engages" with the steel disc , as a slip clutch
this m/c has been stored in the studio over the years, so no extreme hot or cold. So as no sign of rust , either the plastic shrinks with age or there is stress set into the plastic when formed around the steel and then the plastic weakens with age
Shaving back the steel allowed resetting the ends, a ring of heatshrink around and carefully heated the heatshrink , only, with soldering iron and glued around whole disc/plastic join. Looks a goood , non-jamming and non-oval repair, will see tomorrow, get it back in the deck
Back to the original problem. Laces and unlaces and plays , FF and REW ok. Looks as though the gear broke over time in storage.
The PCM legend flickers randomly and no sound out, no 6ch VU bargraph display , the original problem. Not tried playing a video tape but I expect no picture either. Other than cleaning all tape running surfaces , anything to check for? does not like PLAY+REW drops out. Reminds me of Beta video failure mode, usually slip clutch problems ISTR
I decided to inspect under the deck, easier than first seemed. Now with a better viewing angle for the heads, one of the 3 is broken.
Video 8 head disc, 3 heads but 4 coils marked
A and B on opposite sides of drum
head FE ,and combined head JOG on same mount as B
no pic found on gogle pics
Trying to determine what a doubled-up head loks like , as before demounting the drum I thought it was 3 coils , with one of them the ferrite gap section broken away , but now inside looks like 2 purpose made separated ferrite end forms, ie not fractured. On Sony EV S700 , from SM, FE is probably Flying Erase. I may attempt a pic of what is here
a close-up pic of Sony EV S700 , video 8 JOG, doubled up head
ignore the well in the aluminium, dark enamelled coil on left, clear enamel on right and two dark glass/ceramic linear features with a large gap, compared to normal gap, as designed or part broken away?
As rounded edges, I would say as designed
is the only pic I can find
one broken and one good dual head.
I see now that the central ceramic pillar of only about 0.2mm width that has a few shared turns of dark and clear wire would have had an extension out and then 2 now missing gaps, no wonder they break , there is hardly any meat to that sort of construction
arranging the light and viewing a bit better, one gap and its ferrite pillar is still present. I think I'll try grinding up some ferrite with epoxy , prefit a piece of mica for the gap and add a blob of ferrite plus epoxy. If its possible to retrieve something then probably better than nothing, if a few hours of lowest frequencies of the most significant audio can be retrieved. Will try on insignificant tapes first and wait for something to turn up on ebay and keep this in reserve for the no-turn-up situation (assuming anything comes out that is)
There are 4/5 head coils, 4 pairs of wires going down tothe rotary coupler coils but 3 lumps of brass with heads on, one head with 2/3 coils on it. I think the complex one was 3 coils but 2 gaps
The service manual shows 4 heads
What is the connection between Sony & Pioneer?
Pioneer VE D70
and sony EV s700
certainly look the same, different r/c , not found a Pioneer SM yet
CD units Sony PD-C7 and Pioneer CK-W700 of mid 1980s also the same
> From what I heard Sony made ALL 8 mm. decks no matter what they're branded. Makes sense, nobody else wanted to tool up for it so they just bought them. Also there was just a bit of a format war with the rightfully dead VHS-C.
> This is probably good for you because that means that most of them will interchange. Another thing is that it is not like VHS, you can put a head disk in there that does not have the trick head and it will work fine, as opposed to VHS where one set is dedicated to one speed and the other for the other. Actually it CAN be done on VHS by just making the head switch stay in EP and adjusting the PG shifters, if they are DA. If they are discrete heads you are UTCWAP.
> Anyway, if you do find a head that is otherwise compatible without the trick head all that will happen is you won't have a clear still frame or search in SP mode. Big deal. I think mostly all the video tape repairs now should concentrate on just getting it to work and dub the content over to DVD or rip it to a harddrive. Screw the trick head.
> I did a quick search and it seem Asti doesn't have it, I found one online used for $75. Really you could probably find a camcorder somewhere cheap, with maybe something wrong with the camera part. Unlike VHS-C the heads are the same as in the decks. Actually I think I have one kicking around somewhere with a defective auto-focus circuit, but I am not so sure I want to part with it.
Red mark on the drum , not green , SM does not explain the difference
At the moment i'm setting up a "right" pentaprism stereo microscope,or rather an angle-adjustable mount for the video drum, to try fitting a bit of ferrite. Broke a pot core by squashing , so have a goodly collection of ferrite bits of about the right size
I'm always amazed there is enough signal to come off VCR magnetic tape, via little heads and rotary transformer to produce anything of any use
went back to plan A as my temporary glue method failed to fransfer bits of ferrite of sort of dimensions .5x,.3x.3mm.
Epoxy loaded with ferrite grinding dust, seemed to make a good bridge bwetween gap and small coil, probably surface tension helping.
I will leave to cure until tomorrow but anyone know of an out of machine rudimentary test for head functionality? Heads on their own , disconnected from rotary coil, rotating magnet? just so can try on each head ,in turn, to compare responses
Reassembling the deck the mirror plate was out of position due to unsupported side sections can easily move,
the mirror plate must be loosely pivotted. A 1mm rod inserted down the left side just in front of the
mirror plate will delatch the cassette carrier and allow it to rise from internal spring energy, ie no power to deck,
to check for correct posistioning of cassette carrier.
To inspect underside of the mechanical deck 2/3 large sc on the inside top bracing plate and mark and disconnect cons,
comes away quite easily. Then 2 pcb clips and pcb will slide off the other 2 clips.
SM says timing belt and it is actually a toothed belt to the FF/REW drive surprisingly. But small band from motor
although working the drive was off the pulley and acting ove rthe boss unde rthe pulley and band probably stretched a bit
Now failure of REW
Deck has a white lug at the front that sits in a recess of the front panel.
Too much brake on take up side perhaps, on REW
Made a test tape by cutting into the casing the mid section viewable through the
front panel , with flat "letter box" removed, leaving the 2 to sprung pivots
for the spools.
Can then use a needlepoint to assist the relevant over-braked spool , to allow REW or FF
and as confirmation, can also watch central drive cog.
I've never, in earnest , removed an individual head and fixed it to another drum. If enough of the ferrite was remaining on the original I would clamp the drum to a stout plate and manouevre a vinyl record stylus tip adjascent to the gap in the ferrite and fix in place and another such stylus on the other side, before undoing the screw holding the brass in place. Then gently move the replacement head into position and somehow clamp the brass to the drum before tightening the screw
my repair did not work , so whatever turns up on ebay as "needs repair" or even working order
Sony EVS700 VCR, head drum UK / USA compatibility?
Would an electronically dead USA machine be a suitable "parts mule" for a UK machine as far as the video head drum is concerned? NTSC v PAL make a difference to drum or heads physical dimensions or head posistioning etc ?
Required to retrieve sound from tape archives when the PCM , 6 channel audio option ,was used instead of for video use
Unfortunately the desoldering part of drum removal is more awkward than any of the video heads I've removed over the years, so probably the whole machine shipped over
Its last resort time, so a matter of can only give it a go or the tapes remain in archive until a UK machine turns up, if ever
Sony KV2216 tv
Inability to tune in stations and erratic functionig of tuner buttons
One of the program buttons was effectively permanent on and replaced
all similar buttons with switches with a positive click action.
Sony RDR GX210, 2005
Display says LOADING then goes to OPEN
Won't read CDs either, red glow of laser can be seen
one VR on PS is on overlay but not used on this model, P1M01, 2.5V
No HV retaine don mains caps after sw off
Hot side switcher driver Fairchild DF41/ DM0565R ?
Fan runs at 2 speed 6.4V low
Sony RW DVD optical deck , the only part ident seems to be a paper QR label
marked on the optics carrier.
and under the graphic
Googling any of those numbers with or without the extra 0 produces nothing.
No other part mumbers seen . Is the QR world totally separate from the www
world? or just no info out there , full stop.
Is that extra 0 like lopping off the (UK anyway) supermarket shelf label bar
code number to match the actual printed on the product bar code label number
4B WXJ , ZIF on sleedge, lever up , hinged on pcb side .
Kyocera conn at rear of deck processing board jeweller's screwdriver to side at edge
of ribbon , hinge side , pcb end
Video board ZIF, pull up, retaining towards ribbon
MSDH W006A 22B521 on ribbon
2 SMD presets commoned to ground
Service manual out there for this, but nothing on the optical deck ,
DR-02 IPT-2-1, no useful info out there
Sony part number comes up as nothing
2 SMD presets available to twiddle near the lasers
originally 302R and 470R readings
Decreased both, so readings of 254R and 391 , no disc reading
Turned up to 329R and 540R
Now reads all CDs and 4 out of 5 times reads DVDs, I've not (wasted) a DVD-R
on this yet.
Continue twiddling a bit ?
Would be nice to know what these 2 presets do, is there any convention in
relation to closeness to
the 2 lasers as to which one is which. Both are returned to ground.
Now know where these presets are its possible to leave the ribbons in place, move
the sledge to centre posistion and get a blade to the presets
Tried presets at 303& 478 (nearest platter) then 332/540 , no change
Beware of foil ribbon under main drive board LCN 201
FM to NEC C330GC pin 120
RF+ to pin 17, trst point TP228
BA3258 p4 1.5V, p5 3.3V
Unclip Cirrus video pcb end of ribbon as more robust con.
Pull drawer motor ribbon , only, to shift deck pcb to gain access to the presets
Set to 371 and 593R and trying CDs and DVDs read 12 out of 12 in succession
RF signal less than 2mV
R/C RMT D215P
CD-R not recognised
DVD-R recognised and recorded a test disc to of static image and stereo 1KHz mainly , then ea ch separate and then
10K, 100 and 10Hz
UOSD 2000 remote
Battery contacts need re-soldering with wire reinforcement on
the pcb and hot-melt glue filling the gap between the contacts
and the main IC to hold in place.
TV remote (zapper)faults and repairs.
1;resolder battery contacts on PCB and physically
2;IR LED solder joints,sometimes IR LED failure (solder any old
red LED in parallel to check).
3;Broken tracks from being sat on/knelt on
4;failed ceramic resonator,check
by replacing with any old resonator before
aquiring the correct frequency.
5;Gradual failure of the most used buttons-worn
conductive pads-see hints and tips files.
6;Sudden loss of some functions due to break in track in
the keyboard multiplexing.
7;Very rarely is the fault in the main IC.
8;The most common problem is spilt liquid corroding the contacts,
through-board jumpers or pcb tracks.
Saitech 165H (Kasparov) Chess computer
Failure to acknowledge presence of pieces on 2
of the rows.
Liquid had got between the playing surface and the control
pad area where the silver printed signal lines pass through
and corroded the tracks.Hard wire with fine copper
wire between the line area (use silver loaded paint to contact)
and the main pcb.Place tape over the gap to avoid any future
trouble cleaning with an over damp cloth etc.
Burnt and burst SMPS supervisor on this otherwise no-name and little chance Chinese thing. As laser and
motor/s in the DVD deck and heatsink bonded to a video chip a reasonable power draw.
But the SMPS IC , in a conventional 8 pin DIP package also contains the Tx driver transistor. So air gap under
the IC and reverse side of pcb enlarged pad of pcb-copper as some sort of vain attempt at heatsink. Thermal
connection to this is via just pins 4 and 5 combined in normal soldering , connected to the Tx. There is a hole
through the pcb as though for some sort of clip to go perhaps top and bottom of the IC and bolted through to
this under pad but no sign of anything having been fitted there, no deformation of the tinning over the pad
around the hole. Turning over the unit and shaking over paper meant the blasted away top part , mainly, of the
IC emerged but no clip at all.
Is it possible to have combined supervisor and driver transistor in an 8pin DIP package with no heatsink tab
provision? or are there combined SMPS TX with internal driver transistor buried in the base?
Also ps supply provision marked on pcb ac 85V - 250V
IC logi i in a reversed D
marked in 3 lines ICE 2?26 maybe more numbers, ? maybe a 4
logo then 0 and more
XK4.. probably more numbers
p1 0.1uF to 0V
p2 to optocoupler
p3 over current sense
p4-p5 to Tx low side
p7 supply <25V
Its the first time I've come across such a situation, still does not look right. For something like a set-top box
with no motors or laser and a graphics processor that does not require a heatsink, I could see it was possible
courtesy of alldatasheet with "start with" option
ICE2 starting it seems its like a
ICE2A265 with a different logo, same pinning on p4/5/6 at least
large datasheet , will have to read it when I get some time
Perhaps another piece of hitech saved from the dumpster
i in reverse D ,logo on
as infineon, perhaps older or IC only logo (don't know German) and the other pinnings agree with Infineon
datasheet. Will get a few from RS , but will add an underpad-soldered copper 20mm pcb fuse clip and white
goo around the IC, assuming it will still oscillate etc
Replaced .33R current sense r and replaced mains fuse with 630mA , not 2A , as 240V use
Tried replacement IC driver with vaned h/s over, but may have been cause of
non functioning, although no shorting to pins etc.
1V on Vcc
As datasheet made special mention of circuit layout.
Replaced with a new one and no h/s and fired up fine
12.6V on supply pin and display , LEds and kick of CD , so removed the 18R and checked out properly, no
alarming heating of the driver IC
Telephone number - the same number as it has been since 1988
but email is now the preferred method of contact so number deliberately not placed here.
I devote time each day to replying to emails.
(obscure/obsolete components,second hand test equipment,
Postal: 66 Ivy Rd,St Denys,Southampton,England SO17 2JN
There is no point in contacting me about any of the above, the
repair job may have been done 15 years ago .
I cannot clarify or enlarge on any of the above.
A reserve email account is diverse9(commercial at)fastmail.fm.
Please make emails plain text only , no more than 5KByte or 500 words.
Anyone sending larger texts or attachments such as digital signatures, pictures etc will have
them automatically deleted on the server. I will be totally unaware of this - sorry, again
blame the spammers.